### v1

# On optimal solutions of classical and sliced Wasserstein GANs with non-Gaussian data

Abstract
The generative adversarial network (GAN) aims to approximate an unknown distribution via a parameterized neural network (NN). While GANs have been widely applied in reinforcement and semisupervised learning as well as computer vision tasks, selecting their parameters often needs an exhaustive search and only a few selection methods can be proved to be theoretically optimal. One of the most promising GAN variants is the Wasserstein GAN (WGAN). Prior work on optimal parameters for WGAN is limited to the linear-quadratic-Gaussian (LQG) setting, where the NN is linear and the data is Gaussian. In this paper, we focus on the characterization of optimal WGAN parameters beyond the LQG setting. We derive closed-form optimal parameters for one-dimensional WGANs when the NN has non-linear activation functions and the data is non-Gaussian. To extend this to high-dimensional WGANs, we adopt the sliced Wasserstein framework and replace the constraint on marginal distributions of the randomly projected data by a constraint on the joint distribution of the original (unprojected) data. We show that the linear generator can be asymptotically optimal for sliced WGAN with non-Gaussian data. Empirical studies show that our closed-form WGAN parameters have good convergence behavior with data under both Gaussian and Laplace distributions. Also, compared to the r principal component analysis (r-PCA) solution, our proposed solution for sliced WGAN can achieve the same performance while requiring less computational resources.
I Introduction
Generative adversarial networks (GANs) are machine learning frameworks put forth by Goodfellow et al. [1]. A GAN aims to learn an unknown distribution from training data via two competing components, namely the generator and the discriminator. The former tries to mimic the distribution of training data while the latter discriminates between true data and generated data. This framework formulates a minimax problem, where the discriminator aims to maximize the probability of differentiating true data from generated ones, while the generator seeks to minimize this probability [1]. Besides computer vision tasks, applications of GANs to communication systems have also received a lot of attentions. For example, GANs have been applied to autonomous wireless channel modeling [2][3] and mmWave MIMO channel estimation [4].
Traditionally, both the generator and discriminator in GAN are approximated by neural networks (NNs) [1, 5]. Without the NN restrictions and under an optimal unconstrained discriminator, the minimax problem associated with a GAN becomes the minimization of the Jensen-Shannon divergence (JSD) between the distributions of true and generated data. However, due to the nature of this minimax game, the GAN suffers from several problems including vanishing gradient and mode collapse. Many variants of GAN have been proposed to solve these problems, and one of the most promising variants is the Wasserstein GAN (WGAN) [6] , which replaces the JSD by the Wasserstein distance from the optimal transport theory [7]. The WGAN is differentiable with respect to the generator parameters almost everywhere, which benefits the convergence of stochastic gradient descent (SGD) widely adopted for training NNs [6].
Despite the substantial success of applying WGAN to learn distributions in real applications, there are only a few GAN parameter selection algorithms proved to be theoretically optimal [8, 9], which limits the development of GAN beyond heuristic methods in [1, 5]. This lack of rigorous analysis also restricts the evaluation of GANs’ performance to subjective terms. One exception is [8] where Feizi et al. attempted to theoretically understand WGANs on a simple linear quadratic Gaussian (LQG) setting. In this benchmark setting, the synthetic data is generated by a Gaussian distribution, the generator NN is restricted to be linear, and the loss function is quadratic. It is shown in [8, Theorem 1] that for this simple setting, the optimal GAN solution happens to be the principal component analysis (PCA) solution. Regularized versions of GANs are also considered and analyzed [5]. Optimal WGAN solutions under LQG settings, with additional entropic and Sinkhorn regularizers, are also studied [9].
In this paper, we aim to analytically solve WGANs beyond the LQG setting. As described in Sec. II, our setting allows non-Gaussian data distribution and non-linear generators constructed from sigmoid and ReLU functions, and is therefore more general than [8, 9, 10]. Also, we attempt to exactly solve WGANs rather than their regularized versions as in [9]. All of these make our problem much more challenging—after all, even for the inner discriminator problem, which is an optimal transport problem, the solution in most cases is numerically approximated but not analytically characterized [7]. To overcome the challenge, we first focus on one-dimensional data and generator in Sec. III-A, where we provide closed-form solutions for optimal generators in one-dimensional WGANs defined in Sec. II. For high-dimensional WGANs, we adopt the sliced WGAN framework and define a revised sliced WGAN with unprojected data distribution in Sec. II. We provide closed-form solutions for asymptotically optimal generators in high-dimensional sliced WGAN in III-B. The proofs are presented in Sec. IV, where we leverage on results in [11] to solve the inner discriminator problem in closed-form, which greatly simplifies the necessary conditions of optimal WGAN parameters. Moreover, our closed-form solutions do not need any training for the discriminator as [12] and hence provide additional benefit for training WGAN with a decentralized system [13]. Empirical studies in Sec. V show that our closed-form one-dimensional WGAN parameters have good convergence behavior with synthetic data under both Gaussian and Laplace distributions. Empirical studies also show that our closed-form high-dimesional sliced WGAN can acheive the same performance as r-PCA solutions, while requiring fewer computational resources.
Throughout the paper, we adopt the following notational conventions. The operator $\mathrm{diag}(\cdot):\mathbb{R}^{d\times d}\to\mathbb{R}^{d}$ extracts the diagonal entries of a $d\times d$ matrix and returns them as a $d$-dimensional vector. We let $\mathbf{1}\in\mathbb{R}^{d}$ denote the vector whose components are all equal to one, and $e_{i}\in\mathbb{R}^{d}$ denote the $i$-th standard unit vector, with its $i$-th component equal to $1$ and all others equal to 0.
II Problem Formulation
For the WGAN considered in this paper, the generator function is denoted by $G_{\Theta}(.)$, where $\Theta$ is the generator NN’s parameter (weights). Let $q\in\{1,2\}$ represent the order of the Wasserstein distance. In a $q^{\mathrm{th}}$-order WGAN setting, one aims to solve
$\min_{\Theta}\left(\inf_{\pi\in\Pi(\mu,\nu^{\Theta})}\mathbb{E}_{\pi}[\|X-G_{\Theta}(Z)\|^{q}]\right)^{1/q}$ (1)
for an optimal parameter $\Theta$, where $\|\cdot\|$ denotes the $\ell_{2}$-norm in $\mathbb{R}^{d}$, $\mu$ and $\nu^{\Theta}$ are probability measures on $\mathbb{R}^{d}$, generated by the data $X$ and the generator output $G_{\Theta}(Z)$ for a given $\Theta$ and Gaussian input $Z$, respectively. Also, $\Pi(\mu,\nu^{\Theta})$ is the set of probability measures on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ with marginal constraints $\mu$ and $\nu^{\Theta}$, respectively, satisfying $\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{q}d\mu(x)d\nu^{\Theta}(y)<\infty.$ The WGAN problem described in (1) can be equivalently written as $\min_{\Theta}(\mathbb{E}_{\mu}[P(X,\Theta)])^{1/q}$ where the inner-discriminator problem is defined as
$\mathbb{E}_{\mu}[P(X,\Theta)]:=\inf_{\pi\in\Pi(\mu,\nu^{\Theta})}\mathbb{E}_{\pi}[\|X-Y\|^{q}].$ (2)
Note that (2) belongs to the family of the optimal transport problems with $q^{\mathrm{th}}$-order Wasserstein distance, $q\in\{1,2\}$ [7, Proposition 2.2]. In this paper, we consider the following activation functions: 1) the linear function; 2) the rectified linear unit (ReLU) function $\max(0;z)$; and 3) the sigmoid function $1/(1+\exp(-z)).$
Following [1][8], (1) is a population WGAN problem, where $\Theta$ can be optimized with the true data distribution $\mu$, and studying the theoretical population WGAN can lead to practical design with large enough empirical data. In the next Sec. III-A, we first theoretically solve the population WGAN and obtain the closed-form solutions under $d=1$.
Note that for WGAN (1), the closed-form solution was only obtained in the LQG setting [8]. In the LQG setting, the $d$-dimensional synthetic data is generated by Gaussian distribution, i.e., $X\sim\mathcal{N}(0,\mathbf{K})$, the generator is restricted to be a linear generator of the form
$\Theta Z,\quad\Theta\in\mathbb{R}^{d\times r},$ (3)
with $Z\sim\mathcal{N}(0,\mathbf{I}_{r})$ a Gaussian vector, and the loss function is quadratic (i.e., $q=2$).
Beyond the LQG setting, we show how to apply our $d=1$ results for (1) to cope with higher dimensional data $d>1$ by adopting the sliced Wasserstein distance, which offers a computationally efficient alternative to the WGAN [14, 15, 16]. Let $\Omega=\{\omega\in\mathbb{R}^{d}:~\|\omega\|=1\}$ contains all the directions in $\mathbb{R}^{d}$. In sliced WGAN, we project both the data and the generator’s output onto $\mathbb{R}$ through a random direction $\omega\in\Omega$ and compute the Wasserstein distance between the projected 1-dimensional distributions $\mu_{\omega}$ and $\nu^{\Theta}_{\omega}$, i.e., $\omega^{T}X$ and $\omega^{T}G_{\Theta}(Z)$ respectively follow distributions $\mu_{\omega}$ and $\nu^{\Theta}_{\omega}$. By replacing the Wasserstein distance of inner-discriminator problem (2) with sliced one,
$\mathbb{E}_{\mu_{\omega}}[P(X,\Theta)]:=\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega,$ (4)
the $q^{\mathrm{th}}$-order sliced (population) WGAN is
$\min_{\Theta}\mathbb{E}_{\mu_{\omega}}[P(X,\Theta)]^{\frac{1}{q}}$
$=\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\right)^{\frac{1}{q}}.$ (5)
Moreover, though widely adopted in [14, 15, 16], the inner-discriminator problem (4) for the sliced WGAN (II) is optimized under randomly projected data marginal constraint $\mu_{\omega}$. Compared with (2), it is reasonable to change the marginal constraint on data of $\pi$ from projected $\mu_{\omega}$ to unprojected $\mu$,
$\pi\in\Pi(\mu,\nu^{\Theta}_{\omega}),$
and the new sliced WGAN problem is
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\right)^{\frac{1}{q}}.$ (6)
Note that for original (II) and the new (6) with unprojected $\mu$, the optimal values are both non-negative. Thus, we say that a generator $G_{\Theta}(Z)$ is asymptotically optimal for (II) if the corresponding value goes to zero as $d\rightarrow\infty$. The same definition applies to (6). In Sec. III-B, we identify the asymptotically optimal generator for both original and unprojected sliced WGAN problems (II) and (6).
III Main Results
In Sec. III-A, we analytically solve WGAN (1) with the non-linear generator and non-Gaussian data under $d=1$ when $q=2$. With large $d$ in Sec. III-B, we show that the linear generators can asymptotically optimal for both original and unprojected sliced WGANs (II) and (6) by solving corresponding upper-bounds.
III-A Results for WGAN under one dimension
First, we consider the quadratic case $q=2$ with a non-linear generator
$G_{\Theta}(Z)=\theta_{1}+\theta_{2}h(Z),$ (7)
where $h:\mathbb{R}\to\mathbb{R}$ and $Z\sim N(0,1)$, also $(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}$ are parameters of the generator NN to be selected. Let $\Psi$ denote the cumulative distribution function (CDF) of $h(Z)$, for any continuous data distribution $\mu$, our closed-form WGAN parameters are as follows:
Theorem 1.
Assume CDF $F_{\mu}$ of $\mu$ and CDF $\Psi$ of $h(Z)$ in (7) are continuous and strictly increasing, and variance $\operatorname{Var}(h(Z))>0$. If
$\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))+\Psi^{-1}(1-F_{\mu}(X))\right)\geq 0,$ (8)
the population WGAN (1) with $q=2,d=1$ has a unique minimizer for $(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}$ as
$\begin{split}\theta^{*}_{2}&=\frac{\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\geq 0,\\ \theta^{*}_{1}&=\mathbb{E}_{\mu}[X]-\theta^{*}_{2}\mathbb{E}_{g}\left[h(Z)\right];\end{split}$ (9)
if (8) is not met, $(\theta_{1}^{*},\theta_{2}^{*})$ is given by replacing $\theta^{*}_{2}$ in (9) by
$\begin{split}\theta^{*}_{2}&=\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\leq 0,\end{split}$ (10)
where $\mathbb{E}_{g}$ is taking expectation over Gaussian $Z\sim\mathcal{N}(0,1)$.
Proof:
To solve the inner discriminator problem (2), we break (1) down into two sub-problems depending on the sign of $\theta_{2}$, i.e., $\min_{\theta_{1},\theta_{2}\in\mathbb{R}}\mathbb{E}_{\mu}[P(X,\theta_{1},\theta_{2})]$ equals to
$\min\!\left(\!\min_{\theta_{1}\in\mathbb{R},\theta_{2}\geq 0}\mathbb{E}_{\mu}[P(X,\theta_{1},\theta_{2})],\!\min_{\theta_{1}\in\mathbb{R},\theta_{2}\leq 0}\mathbb{E}_{\mu}[P(X,\theta_{1},\theta_{2})]\!\!\right)$ (11)
The solution of the first sub-problem is indeed (9) while that for second sub-problem is (10). The condition (8) is obtained by comparing the values of the two subproblems. The following proof is given in Sec. IV-A. ∎
Let us now look at some specific cases of $h(z)$. For the sigmoid function $h(z)$, recall that the logit function $\text{logit}(p):=\ln\left(p/(1-p)\right),p\in(0,1)$ which is the inverse function $h^{-1}(z)$. The random variable $h(Z)$ has a logit-normal distribution, i.e. $\text{logit}(h(Z))$ is normally distributed, with CDF
$\Psi(v)=\frac{1}{2}\left(1+\text{erf}\left(\frac{\text{logit}(v)}{\sqrt{2}}\right)\right)\quad\hbox{for}\ v\in(0,1).$
For the ReLU function, the CDF of $h(Z)=\max\{Z,0\}$ is given by
$\Psi(v)=\Phi(v)\cdot 1_{\{v\geq 0\}},$ (12)
where $\Phi$ is the CDF of Gaussian $N(0,1)$. However, now $\Psi$ has a jump from 0 to $1/2$ at $v=0$ and does not meet the setting of Theorem 1 since it is neither continuous nor strictly increasing. We need the following modification.
Theorem 2.
Assume $\mu$ has the setting as in Theorem 1 and $\Psi$ is given by (12). If
$\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)$
$\geq$ $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right),$ (13)
the population WGAN (1) with $q=2,d=1$ has a unique minimizer for $(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}$ as
$\begin{split}\theta^{*}_{2}&=\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)\\ \theta^{*}_{1}&=\mathbb{E}[X]-\theta^{*}_{2}/\sqrt{2\pi};\end{split}$ (14)
if (13) is not met, $(\theta_{1}^{*},\theta_{2}^{*})$ is given by replacing $\theta^{*}_{2}$ in (14) by
$\theta^{*}_{2}=-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right).$ (15)
The proof of Theorem 2 is in Appendix A. For the linear generator $h(Z)=Z$ as [5, eqn. (27)][8][9], one can simplify the results in Theorem 1 and also get alternative closed-form formula as follows.
Corollary 1.
Assume $\mu$ has the setting as in Theorem 1 and consider the linear case $h(Z)=Z$ in (7). Then population WGAN (1) has a unique minimizer for $(\theta_{1},\theta_{2})\in\mathbb{R}\times\mathbb{R}$ as
$\theta^{*}_{1}=\mathbb{E}_{\mu}[X]\quad\hbox{and}\quad\theta^{*}_{2}=\mathbb{E}_{\mu}[X\cdot\Phi^{-1}(F_{\mu}(X))],$ (16)
also equivalently
$\theta^{*}_{2}=\mathbb{E}_{g}[F_{\mu}^{-1}(\Phi(Z))\cdot Z],$ (17)
The proof of Corollary 1 is in Appendix B.
Here we briefly compare our proposed WGAN solutions with the results for the LQG setting in [8, Theorem 1]. Unlike [8], our results can deal with non-Gaussian data distribution. Moreover, we can recover the result in [8] when $q=2,d=1$. Indeed, if $X\sim\mathcal{N}(0,\sigma^{2})$, by looking into (16), we have $\Phi^{-1}(F_{\mu}(X))=\Phi^{-1}(\Phi(X/\sigma))=X/\sigma.$ Plugging this into our solution in (16) shows that $\theta_{1}^{*}=0$ and $\theta_{2}^{*}=\mathbb{E}[X\frac{X}{\sigma}]=\sigma$, coinciding with the result in [8] for $d=r=1$. We emphasize that neither of our and the results in [8] subsume the other as a special case as [8] considers general dimension $d$ while our work is not restricted to Gaussian data distribution. Since WGAN output is already Gaussian $\Theta Z\sim\mathcal{N}(0,\mathbf{K}^{\prime})$, it is shown in [8, Theorem 1] that for this benchmark, the optimal WGAN solution happens to be the $r$-PCA solution of Gaussian $X$. That is, optimal generator matrix $\Theta$ fulfills that $\mathbf{K}^{\prime}=\Theta\Theta^{T}$ is a rank $r$ matrix and $\mathbf{K}^{\prime}$ and $\mathbf{K}$ share the same largest $r$ eigenvalues and the corresponding eigenvectors.
Finally, we present our result for non-linear generators and first-order Wasserstein distance ($q=1$). We have
Corollary 2.
Following the settings in Theorem 1, the minimizer $(\theta^{*}_{1},\theta^{*}_{2})$ of the population WGAN (1) with $q=1,d=1$ meet the following necessary conditions
$\mathbb{E}_{\mu}\!\!\left[\text{sign}\Big{(}\theta^{*}_{1}+\theta^{*}_{2}\Psi^{-1}(F_{\mu}(X))-X\Big{)}\right]=0,$
$\mathbb{E}_{\mu}\!\!\left[\text{sign}\!\Big{(}\theta^{*}_{1}+\theta^{*}_{2}\Psi^{-1}(F_{\mu}(X))-X\Big{)}\Psi^{-1}(F_{\mu}(X))\right]\!\!=\!\!0,$ (18)
when $\theta^{*}_{2}>0$, where $\text{sign}(x)=1,0,-1$ for $x>0$, $x=0$, $x<0$, respectively.
The proof is omitted. Note that when $X\sim\mathcal{N}(\mu,\sigma^{2})$ then $F_{\mu}(X)=\Phi((X-\mu)/\sigma)$, It can be checked that for the linear case as Corollary 1, the optimal $(\theta^{*}_{1},\theta^{*}_{2})=(\mu,\sigma)$ for $q=2$ also meets (18) and is at least a local optimum for $q=1$.
III-B Results for sliced WGAN under high dimension
We first focus on the asymptotic optimal generator of (II) as defined in Sec. II. Our main result shows that the linear generator can be asymptotically optimal for the original sliced WGAN (II) when $q=2$.
Theorem 3.
For the second-order sliced WGAN (II), if data $X$ satisfy the weak dependence condition in [16, Definition 1], the linear generator (3) is asymptotically optimal with the parameters
$\Theta^{*}=\mathbf{U}\Lambda\mathbf{V}^{T}\text{ where }\Lambda\Lambda^{T}=\frac{\Gamma(\frac{d+1}{2})^{2}}{2\Gamma(\frac{d}{2}+1)^{2}}\mathbb{E}_{\mu}[\|X\|^{2}]\mathbf{I},$ (19)
$\mathbf{U},\mathbf{V}^{T}$ are unitary, and $\Gamma(\cdot)$ is gamma function.
For the proof sketch, rather than finding the closed-form solution of the original sliced WGAN (II) directly, we solve its upper-bounds with bounded-error and approximations to overcome the the complicated integration and projection of (II). First, by replacing general generator $G_{\Theta}(Z)$ with linear generator (3), (II) is upper-bounded by
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{q}]d\omega\right)^{\frac{1}{q}}.$ (20)
Motivated by [17] we replace the uniformly distributed $\omega\in\Omega$ in (20) with Gaussian projections $\omega\in\Omega_{G}$, that is, $\omega$ is generated according to a zero-mean Gaussian distribution with covariance matrix $(1/d)\mathbf{I}_{d}$ and aim at solving
$\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{q}]d\omega\right)^{\frac{1}{q}},$ (21)
[…]
Lemma 1.
Let $\hat{X}_{G}\sim\mathcal{N}(0,\tilde{\sigma}^{2}_{x})$ independent of Gaussian vector $\omega$ with finite $\tilde{\sigma}^{2}_{x}=\mathbb{E}_{\mu}[\|X\|^{2}]/d.$ The gap of (21) to
$\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}\Theta Z|^{q}]d\omega\right)^{1/q}$ (22)
is bounded by a function $f^{\mu}(d)=O(d^{-1/8})$, where the marginals of $\pi$ are $\mathcal{N}(0,\tilde{\sigma}^{2}_{x})$ and $\nu^{\Theta}_{\omega}$. Solving (22) with the linear generator (3), the optimal $\Theta$ equals to that of
$\min_{\Theta}\mathbb{E}_{\omega}[|\tilde{\sigma}_{x}-\sqrt{\omega^{T}\Theta\Theta^{T}\omega}|^{q}],q=1,2.$ (23)
Moreover, the optimal $\Theta$ when $q=2$ is the minimizer to
$\frac{\mathrm{Tr}(\Theta\Theta^{T})}{d}+\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}z^{-1/2}\frac{\partial}{\partial z}\left|\mathbf{I}+\frac{2z}{d}\Theta\Theta^{T}\right|^{-1/2}dz;$ (24)
where $\mathrm{Tr}(.)$ is the matrix trace.
As summary to the current steps, the upper-bounds of (II) under $q=2$ are
$\min_{\Theta}\left(\int_{\omega\in\Omega}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{2}]d\omega\right)^{\frac{1}{2}}$
$\leq$ $\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]d\omega\right)^{\frac{1}{2}}$
$\leq$ $\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}\Theta Z|^{2}]d\omega\right)^{1/2}+O(d^{-1/8}).$ (25)
where the optimal $\Theta$ in (III-B) can be solved by (24). Then, the final upper-bound in (III-B) becomes
$\tilde{\sigma}_{x}\sqrt{1-\frac{d\Gamma(\frac{d+1}{2})^{2}}{2\Gamma(\frac{d}{2}+1)^{2}}}+O(d^{-1/8}).$ (26)
Indeed one can show that (26) tends to zero as $d\rightarrow\infty$, and thus linear generator in the second upper-bound of (III-B) is asymptomatically optimal for (II). Thus, in the sense of asymptotic optimality defined in Sec. II, even with the non-linear generator, the value of sliced WGAN under very high dimensional non-Gaussian data cannot be smaller than that from proposed linear generator (19).
To solve the value (26) of the upper bound (III-B), we first apply singular value decomposition (SVD) on parameters $\Theta=\mathbf{U}\Lambda\mathbf{V}^{T}$, and further simplify the objective using the Carlson-R function $R_{-a}(\mathbf{b};\mathbf{z})$ [18], where $\mathbf{b},\mathbf{z}$ are $\mathbb{R}^{d}$ vectors. Let $\mathbf{S}=\Lambda\Lambda^{T}$. Now, the optimization problem becomes
$\min_{\mathbf{S}}\frac{\mathrm{Tr}(\mathbf{S})}{d}-\tilde{\sigma}_{x}\sqrt{2d}\frac{\Gamma({\frac{d+1}{2})}}{\Gamma({\frac{d}{2}+1)}}R_{\frac{1}{2}}(\frac{1}{2}\mathbf{1};\mathrm{diag}(\mathbf{S})).$ (27)
Then, we prove that the obejective of (27) is Schur-convex [19], so the global minimum of (27) is acheived when all the diagnol elements of $\mathbf{S}$ are equal. Thus, we find the optimal parameters (19) of linear generator, and obtain the upper bound value (26). The full proof of solving the upper bound is given in Sec. IV-B.
Remark 1.
From [16], under the weak dependence condition, the elements of data vector $X$ can follow Gaussian process, Bernoulli shift and autoregressive (AR) model. For channel prediction, AR model-based method is widely used by using past channel data without requiring current channel measure via pilot [20].
Remark 2.
As a sanity check, when $d=2$ and $q=2$, (22) with linear generator can also be solved by elliptic integrals. The solution are the same as (19). The proof is given in appendix E.
For the unprojected sliced WGAN (6), we show that the linear generator is also asymptotically optimal when $q=2$.
Theorem 4.
Under assumption of Theorem 3 with $\mu$ vanishing on hypersurfaces in $\mathbb{R}^{d}$, for the unprojected sliced WGAN (6), the linear generator (3) with the same parameters as (19) is asymptotically optimal when $q=2$.
Proof.
We show that (6) is indeed equivalent to (II) when $q=2$ by the following lemma.
Lemma 2.
For any $\omega\in\Omega$, the values of the inner-discriminator probelms of (6) and (II) with the linear generator are the same when $q=2$,
$\inf_{\pi\in\Pi(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]$ (28)
$=$ $\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}].$
We prove that the optimal transport function $\pi^{*}$ of the L.H.S. in (28) equals to that of R.H.S. in (28). First, the L.H.S. in (28) belongs to the general multi-to-one-dimensional optimal transport problems with closed-form solutions identified in [21]. Then, we identify its analytical properties to prove that the optimal transport function of the L.H.S. of (28) is well-defined and equals to that of the R.H.S.. The detailed proof of Lemma 2 is in Sec. IV-C. Finally, we can prove the desired result for (6) following steps to prove that for original sliced WGAN (II) in Theorem 3. ∎
For application, our upcoming simulations in Sec. V shows that the linear generator (19) can be a low-complexity generator with empirical data. For data without weak dependence condition, the state-of-art WGAN solver can be obtained via gradient-flow of stochastic differential equation [22][23], and our optimal can be a low-complexity initialization and has the potential for faster convergence.
IV The Proofs
Here we give proofs for our main theorems.
IV-A Proof for Theorem 1
In the following, we first solve the first sub-problem with $\theta_{2}\in\mathbb{R}_{+}$ in (11). With $d=1$, we recall the following result for the inner-discriminator problem (2) of (1), by taking $\nu^{\Theta}:=P_{G_{\Theta}(z)}$, the measure generated by the generator NN (with parameter $\Theta:=(\theta_{1},\theta_{2})$ in (7)). Let $F_{\mu}$ and $F_{\nu^{\Theta}}$ denote the cumulative distribution functions (CDFs) of the measures $\mu$ and $\nu^{\Theta}$ on $\mathbb{R}$.
Lemma 3 ([11], Theorem 5.1).
Define $t^{\Theta}:\mathbb{R}\to\mathbb{R}\cup\{\infty\}$ by
$t^{\Theta}(x):=\sup\{y\in\mathbb{R}:F_{\nu^{\Theta}}(y)\leq F_{\mu}(x)\}.$ (29)
For $q=1,2$, if $\mu$ has no atom ($\mu$ is a continuous distribution)
$\inf_{\pi\in\Pi(\mu,\nu^{\Theta})}\int_{\mathbb{R}\times\mathbb{R}}|x-y|^{q}d\pi(x,y)=\int_{\mathbb{R}}|x-t^{\Theta}(x)|^{q}d\mu(x),$ (30)
For the inner discriminator problem (2) with $q=2,d=1$, $\mathbb{E}_{\mu}[P(X,\theta_{1},\theta_{2})]$ for given $(\mu,\theta_{1},\theta_{2})$ equals to (30), where $t^{\Theta}(x)$ is defined as in (29). Now we need to find a closed-form $t^{\Theta}(x)$ to continue. From (7), let $\Psi$ denote the CDF of $h(Z)$. If $\Psi$ is continuous and strictly increasing and $\theta_{2}\in\mathbb{R}_{+}$, (29) can be expressed in closed-form as
$t^{\Theta}(x)=\theta_{1}+\theta_{2}\Psi^{-1}(F_{\mu}(x)).$ (31)
To see this, since $\theta_{2}>0$, observe that
$\Psi\bigg{(}\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}\bigg{)}=\mathbb{P}\bigg{(}h(Z)\leq\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}\bigg{)}$ (32)
$=$ $\mathbb{P}\left(G_{\Theta}(Z)\leq t^{\Theta}(x)\right)=F_{\nu^{\Theta}}\left(t^{\Theta}(x)\right)\overset{(a)}{=}F_{\mu}(x),$
where the last equality follows from (29) and that $F_{\nu^{\Theta}}$ is continuous and strictly increasing; here, $F_{\nu^{\Theta}}$ inherits the same properties from $\Psi$ thanks to (7). Though the continuity of $\Psi$ is not needed in Lemma 3 and its Kantorovich equivalence [24, Theorem 2.18], without it equality (a) will become an inequality which harms finding closed-form $\Theta$. Then it follows that $\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}=\Psi^{-1}(F_{\mu}(x))$, which yields (31). On the other hand, if $\theta_{2}=0$, since $G_{\Theta}(Z)\equiv\theta_{1}\in\mathbb{R}$, we have $F_{\nu^{\Theta}}(y)=1_{\{y\geq\theta_{1}\}}$. Plugging this into (29) directly gives $t^{\Theta}(x)=\theta_{1}$ for all $x\in\mathbb{R}$. This particularly shows that (31) is also satisfied for the case $\theta_{2}=0$. Our target sub-problem becomes
$\min_{\theta_{1}\in\mathbb{R},\theta_{2}\geq 0}\int_{\mathbb{R}}|x-t^{\Theta}(x)|^{2}d\mu(x).$ (33)
As will be explained below, the other sub-problem can be solved in the same way.
With closed-form representation (30)(31) for the inner problem (2), WGAN (1) with $q=2,d=1$ can be simplified to be the following stochastic minimization problem
$\min_{\theta_{1}\in\mathbb{R},\theta_{2}\in\mathbb{R}_{+}}\mathbb{E}_{\mu}\left[\left|X-\theta_{1}-\theta_{2}\Psi^{-1}(F_{\mu}(X))\right|^{2}\right].$ (34)
Together with (7), (34) becomes the constrained optimization problem
$\min_{\theta_{1},\theta_{2}\in\mathbb{R}}\quad\!J(\theta_{1},\theta_{2})\!:=\!\int_{\mathbb{R}}\Big{(}\theta_{1}\!+\!\theta_{2}\Psi^{-1}(F_{\mu}(x))\!-\!x\Big{)}^{2}\!F_{\mu}^{\prime}(x)dx$
$\hbox{subject to}\quad g(\theta_{1},\theta_{2}):=-\theta_{2}\leq 0.$ (35)
The corresponding first-order Karush-Kuhn-Tucker (KKT) condition is
$\nabla J(\theta_{1},\theta_{2})+\lambda\nabla g(\theta_{1},\theta_{2})$ $=0,$ (36)
$\lambda g(\theta_{1},\theta_{2})$ $=0,$ (37)
where $\lambda\geq 0$ is the Lagrange multiplier. By direct calculation, (36) becomes
$\int_{\mathbb{R}}\Big{(}\theta_{1}+\theta_{2}\Psi^{-1}(F_{\mu}(x))-x\Big{)}F_{\mu}^{\prime}(x)dx=0,$
$\int_{\mathbb{R}}\Big{(}\theta_{1}+\theta_{2}\Psi^{-1}(F_{\mu}(x))-x\Big{)}\Psi^{-1}(F_{\mu}(x))F_{\mu}^{\prime}(x)dx=\frac{\lambda}{2}.$
Recall $X$ is a random variable with CDF $F_{\mu}$, and the above equalities can be written as
$\theta_{1}+\theta_{2}\mathbb{E}\left[\Psi^{-1}(F_{\mu}(X))\right]=\mathbb{E}[X],$
$\theta_{1}\mathbb{E}\left[\Psi^{-1}(F_{\mu}(X))\right]+\theta_{2}\mathbb{E}\left[(\Psi^{-1}(F_{\mu}(X)))^{2}\right]$
$-\mathbb{E}\left[X\Psi^{-1}(F_{\mu}(X))\right]=\lambda/2.$
Note that $F_{\mu}(X)\sim\operatorname{Uniform}[0,1]$ from [25, Lemma 1], so that the CDF of $\Psi^{-1}(F_{\mu}(X))$ is simply $\Psi$. In other words, $\Psi^{-1}(F_{\mu}(X))$ and $h(Z)$ have identical distribution. The formulas above thus simplify to
$\theta_{1}+\theta_{2}\mathbb{E}\left[h(Z)\right]-\mathbb{E}[X]=0,$
$\theta_{1}\mathbb{E}\left[h(Z)\right]+\theta_{2}\mathbb{E}\left[(h(Z))^{2}\right]-\mathbb{E}\left[X\Psi^{-1}(F_{\mu}(X))\right]=\frac{\lambda}{2}.$
Plugging $\theta_{1}=\mathbb{E}[X]-\theta_{2}\mathbb{E}\left[h(Z)\right]$ from the first equality into the second one, we obtain
$\mathbb{E}[X]\mathbb{E}\left[h(Z)\right]+\theta_{2}\text{Var}(h(Z))-\mathbb{E}\left[X\Psi^{-1}(F_{\mu}(X))\right]-\lambda/2=0.$
Hence, a solution $(\theta_{1},\theta_{2},\lambda)$ to (36) must equivalently satisfy
$\begin{split}\theta_{1}&=\mathbb{E}[X]-\theta_{2}\mathbb{E}\left[h(Z)\right],\\ \lambda/2&=\theta_{2}\text{Var}(h(Z))-\text{Cov}(X,\Psi^{-1}(F_{\mu}(X))).\end{split}$ (38)
To solve the KKT condition (36)-(37) for candidate minimizers, we already know that (36) boils down to (38), while (37) simply implies either $\lambda=0$ or $\theta_{2}=0$. Also, because $\Psi^{-1}$ is strictly increasing and $F_{\mu}$ is nondecreasing, the map $x\mapsto\Psi^{-1}(F_{\mu}(x))$ is nondecreasing. This readily implies
$\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right)\geq 0$ (39)
in (38) since $(x-x^{\prime})(\Psi^{-1}(F_{\mu}(x))-\Psi^{-1}(F_{\mu}(x^{\prime})))\geq 0$ for all $x,x^{\prime}\in\mathbb{R}$. Specifically, by taking an independent but same distribution copy $X^{\prime}$ of $X$,
$0$ $\leq\mathbb{E}\left[(X-X^{\prime})\left(\Psi^{-1}(F_{\mu}(X))-\Psi^{-1}(F_{\mu}(X^{\prime}))\right)\right]$ (40)
$=2\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right).$
We therefore separate the proof into two cases from (39).
Case I: $\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))>0$. If $\theta_{2}=0$, (38) entails $\lambda=-2\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))<0$, which violates the requirement $\lambda\geq 0$ in (36)-(37). If $\lambda=0$, solving (38) yields
$\begin{split}&\theta^{*}_{1}=\mathbb{E}[X]-\frac{\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))}{\operatorname{Var}(h(Z))}\mathbb{E}\left[h(Z)\right],\\ &\theta^{*}_{2}=\frac{\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))}{\operatorname{Var}(h(Z))}\geq 0.\end{split}$ (41)
That is, the KKT condition gives a unique candidate optimizer $(\theta_{1}^{*},\theta_{2}^{*},\lambda^{*})$, with $(\theta_{1}^{*},\theta_{2}^{*})$ as in (41) and $\lambda^{*}=0$. To check the corresponding second-order condition, let $H_{J}$ and $H_{g}$ denote the Hessian matrices of $J(\theta_{1},\theta_{2})$ and $g(\theta_{1},\theta_{2})$, respectively. Clearly, $H_{g}=O_{2\times 2}$ and is all-zero, which implies
$H_{J}+\lambda H_{g}=2\begin{bmatrix}1&\mathbb{E}[h(Z)]\\ \mathbb{E}[h(Z)]&\mathbb{E}[(h(Z))^{2}]\end{bmatrix}$. (42)
As $\det(H_{J}+\lambda H_{g})/4=\mathbb{E}[(h(Z))^{2}]-\mathbb{E}[(h(Z))]^{2}=\text{Var}(h(Z))>0$, $H_{J}+\lambda H_{g}$ is positive definite. Hence, we conclude that $(\theta_{1}^{*},\theta_{2}^{*})$ in (41) is the unique minimizer of (35).
Case II: $\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))=0$. As (38) entails $\lambda/2=\theta_{2}\text{Var}(h(Z))$, we have $\lambda=\theta_{2}=0$ (by recalling that either $\lambda=0$ or $\theta_{2}=0$). That is, the KKT condition gives a unique candidate optimizer $(\theta_{1}^{*},\theta_{2}^{*},\lambda^{*})=(\mathbb{E}[X],0,0)$. Since $H_{J}+\lambda H_{g}$ is again given by (IV-C), which is positive definite, we conclude that $(\theta_{1}^{*},\theta_{2}^{*})=(\mathbb{E}[X],0)$ is the unique minimizer of (35).
As a remark, another way to prove is to show $H_{J}$ as (IV-C) is positive definite in both cases first, and then $J(\theta_{1},\theta_{2})$ is convex. Since both the proof need to check the Hessian matrix, there is no simpler way to prove.
Lemma 4.
Suppose that $\mu$ has no atom, the cdf $\Psi$ of $h(Z)$ is continuous and strictly increasing, and $\operatorname{Var}(h(Z))>0$. The sub-problem (33) has a unique minimizer $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$, defined as in (9). Moreover, the minimum value is
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)=\operatorname{Var}(X)-\frac{\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))}{\operatorname{Var}(h(Z))}.$ (43)
Proof.
As (33) is equivalent to (35), Theorem 1 directly shows that $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$ as in (9) is the unique minimizer of (33). Thanks to (31) and (9),
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$
$=\int_{\mathbb{R}}\left(x-\theta_{1}^{*}-\theta^{*}_{2}\Psi^{-1}(F_{\mu}(x))\right)^{2}d\mu(x)$
$=\int_{\mathbb{R}}\left(x-\mathbb{E}[X]-\theta^{*}_{2}\Big{(}\Psi^{-1}(F_{\mu}(x))-\mathbb{E}[h(Z)]\Big{)}\right)^{2}d\mu(x)$
$=\int_{\mathbb{R}}(x-\mathbb{E}[X])^{2}d\mu(x)$
$\hskip 14.45377pt-2\theta^{*}_{2}\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))-\mathbb{E}[h(Z)]\Big{)}d\mu(x)$
$\hskip 50.58878pt+(\theta^{*}_{2})^{2}\int_{\mathbb{R}}\Big{(}\Psi^{-1}(F_{\mu}(x))-\mathbb{E}[h(Z)]\Big{)}^{2}d\mu(x)$
$=\operatorname{Var}(X)-2\theta^{*}_{2}\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))+(\theta^{*}_{2})^{2}\operatorname{Var}(h(Z))$
$=\operatorname{Var}(X)-\frac{\operatorname{Cov}(X,\Psi^{-1}(F_{\mu}(X)))^{2}}{\operatorname{Var}(h(Z))},$
where the fourth equality follows from that $\Psi^{-1}(F_{\mu}(X))$ and $h(Z)$ have identical distribution and the last equality is due to the definition of $\theta^{*}_{2}$ in (9). ∎
Lemma 5.
Suppose that $\mu$ has no atom, the cdf $\Psi$ of $h(Z)$ is continuous and strictly increasing, and $\operatorname{Var}(h(Z))>0$. The sub-problem
$\min_{\theta_{1}\in\mathbb{R},\theta_{2}\leq 0}\int_{\mathbb{R}}|x-t^{\Theta}(x)|^{2}d\mu(x)$ (44)
has a unique minimizer $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$, defined by
$\begin{split}\theta^{*}_{1}&=\mathbb{E}[X]-\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\mathbb{E}\left[h(Z)\right],\\ \theta^{*}_{2}&=\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\leq 0.\end{split}$ (45)
Moreover, the minimum value is
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)=\operatorname{Var}(X)-\frac{\operatorname{Cov}(X,\Psi^{-1}(1-F_{\mu}(X)))^{2}}{\operatorname{Var}(h(Z))}.$ (46)
Proof.
Set $\eta:=-\theta_{2}$ and let $\Psi_{0}$ denote the cdf of h(Z)-h(Z). Given $x\in\mathbb{R}$, take $y=\Psi^{-1}_{0}(F_{\mu}(x))$. Observe that $F_{\mu}(x)=\Psi_{0}(y)=\mathbb{P}(-h(Z)\leq y)=\mathbb{P}(h(Z)\geq-y)=1-\Psi(-y)$, which implies $y=-\Psi^{-1}(1-F_{\mu}(x))$. That is,
$\Psi^{-1}_{0}(F_{\mu}(x))=-\Psi^{-1}(1-F_{\mu}(x))\quad\forall x\in\mathbb{R}.$ (47)
Now, since we can rewrite $G_{\Theta}$ in (7) as $G_{\Theta}=\theta_{1}+\eta\cdot(-h(Z))$, the sub-problem (44) can be solved by Proposition 1, with $(\theta_{2},h(Z),\Psi)$ therein replaced by $(\eta,-h(Z),\Psi_{0})$. The unique minimizer $(\theta^{*}_{1},\eta^{*})$ is then given by
$\begin{split}\eta^{*}&=\frac{\operatorname{Cov}\left(X,\Psi^{-1}_{0}(F_{\mu}(X))\right)}{\operatorname{Var}(-h(Z))}=-\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\\ \theta^{*}_{1}&=\mathbb{E}[X]-\eta^{*}\mathbb{E}\left[-h(Z)\right]\\ &=\mathbb{E}[X]-\frac{\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)}{\operatorname{Var}(h(Z))}\mathbb{E}[h(Z)],\end{split}$
where we used (47) to change from $\Phi_{0}$ to $\Psi$. This readily yields (45). By replacing $\Psi$ in (43) with $\Psi_{0}$ and using again (47), we obtain the desired minimum value formula. ∎
Now we can prove Theorem 1. By the linearity of covariances,
$\begin{split}&\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))\right)+\operatorname{Cov}\left(X,\Psi^{-1}(1-F_{\mu}(X))\right)\\ =&\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))+\Psi^{-1}(1-F_{\mu}(X))\right).\end{split}$
If (8) holds, the above implies that the value in (43) is smaller or equal to that in (46). Hence, the minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should be the one associated with the value (43), which is given by (9). On the other hand, if (8) fails, the same argument shows that the value in (46) is smaller than that in (43). The minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should then be the one associated with the value (46), which is given by (45).
IV-B Proof for Theorem 3
In the beginning, here we provide the proof sketch of Lemma 1 and left the details to Appendix C. First, we assume $X$ is zero mean, since equivalently one can modify $G_{\Theta}(Z)$ in (II) with a bias $G_{\Theta}(Z)+\mathbb{E}[X]$ for non-zero mean $X$. With $q=2$, Wasserstein distance (as RHS of (2)) of order 2 between $\omega^{T}X$ with $\hat{X}_{G}$, averaged over $\omega$, is bounded from [17]. We modify this result to make it valid for $q=1$, and then get desired bounded gap $f^{\mu}(d)$ between (21)(22) for $q=1,2$ as
$f^{\mu}(d):=(C^{\prime}\mathbb{E}_{\mu}[\|X\|^{2}](d^{-5/4}+d^{-7/5}))^{1/2}$ (48)
where $C^{\prime}$ is a constant. With linear generator (3), not only $\omega^{T}\Theta Z$ is Gaussian as $\hat{X}_{G}$ given $\omega$, on the contrary to [16, Theorem 1], but also we prevent gap $f^{\mu}(d)$ depending on $\Theta$.
Next (23) is from modifying Corollary 1 and 2 to (22). Finally, solving (23) with $q=2$ equals to solving
$\tilde{\sigma}^{2}_{x}-\max_{\Theta}\mathbb{E}_{w}[2\tilde{\sigma}_{x}\sqrt{\omega^{T}\Theta\Theta^{T}\omega}-\omega^{T}\Theta\Theta^{T}\omega].$ (49)
For the first term, for Gaussian $w\in\Omega_{G}$, let its quadratic form $U_{w}:=\omega^{T}\Theta\Theta^{T}\omega$ and $U_{w}>0$ almost surely. Then, $\mathbb{E}_{w}[\sqrt{\omega^{T}\Theta\Theta^{T}\omega}]=\mathbb{E}_{U_{w}}[U_{w}U_{w}^{-1/2}]$ and we can prove that
$\mathbb{E}_{w}[\sqrt{\omega^{T}\Theta\Theta^{T}\omega}]=\frac{-1}{\Gamma(1/2)}\int^{\infty}_{0}z^{-1/2}\frac{\partial}{\partial z}M_{U_{w}}(-z)dz$ (50)
where $M_{U_{w}}(z):=\mathbb{E}[e^{zU_{w}}]$ is the moment generating function of $U_{w}$. For the second term it is easy to see that
$\mathbb{E}_{w}[\omega^{T}\Theta\Theta^{T}\omega]=\mathrm{Tr}(\mathbb{E}_{w}[\omega\omega^{T}]\Theta\Theta^{T})=\mathrm{Tr}(\Theta\Theta^{T})/d.$ (51)
Combining these results reach (24).
Continuing the proof for Theorem 3, we first aim on solving (24). Though the derivative with respect to $z$ in (24) can be easily solved by matrix calculus [26], directly integration over $z$ is still too complicated for turn it to the closed-form (27). To solve this, applying SVD on the parameters of linear generator $\Theta=\mathbf{U}\Lambda\mathbf{V}^{T}\in\mathbb{R}^{d\times r}$ then $\Theta\Theta^{T}$ can be eigen-decomposed as
$\Theta\Theta^{T}=\mathbf{U}\Lambda\Lambda^{T}\mathbf{U}^{T}=\mathbf{U}\mathbf{S}\mathbf{U}^{T},$
where $\mathbf{U}\in\mathbb{R}^{d\times d}$ is eigenvectors of $\Theta\Theta^{T}$, $\mathbf{V}\in\mathbb{R}^{r\times r}$ is eigenvectors of $\Theta^{T}\Theta$, $\Lambda\in\mathbb{R}^{d\times r}$ and $\mathbf{S}$ is a diagonal matrix. Now diagonal $\mathbf{S}=\Lambda\Lambda^{T}\in\mathbb{R}^{d\times d}$ has each diagonal element $S_{ii}\geq 0$. Substituting $\Theta\Theta^{T}=\mathbf{U}\mathbf{S}\mathbf{U}^{T}$ into (24), it becomes
$\frac{\sum_{i=1}^{d}S_{ii}}{d}+\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}z^{-1/2}\frac{\partial}{\partial z}(\prod_{i=1}^{d}(1+\frac{2z}{d}S_{ii})^{-1/2})dz$ (52)
$=\frac{\sum_{i=1}^{d}S_{ii}}{d}-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}(\prod_{i=1}^{d}(1+\frac{2z}{d}S_{ii})^{-3/2})$
$(\sum_{i=1}^{d}\frac{2}{d}S_{ii}\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj}))dz.$ (53)
To solve the optimal elements of the diagonal matrix $\mathbf{S}$ in (53), we focus on simplifying the integral part of it using the Carlson-R function and turn it to objectives of (27). Details are as follows, in (53),
$-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}(\prod_{i=1}^{d}(1+\frac{2z}{d}S_{ii})^{-3/2})$
$(\sum_{i=1}^{d}\frac{2}{d}S_{ii}\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj}))dz$
$=$ $-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\sum_{i=1}^{d}\int^{\infty}_{0}\frac{1}{d}z^{-1/2}S_{ii}(1+\frac{2z}{d}S_{ii})^{-3/2}$
$\prod_{j=1,j\neq i}^{d}(1+\frac{2z}{d}S_{jj})^{-1/2}dz.$
$\overset{(a)}{=}$ $-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\sum_{i=1}^{d}\frac{1}{d}B(\frac{1}{2},\frac{d+1}{2})S_{ii}R_{-\frac{1}{2}}\left(\mathbf{b}^{(i)};\mathrm{diag}\left(\frac{2}{d}\mathbf{S}\right)\right)$
The step (a) follows [18, (8.3-1)] where $B(\cdot,\cdot)$ is beta function, $\mathbf{b}^{(i)}\in\mathbb{R}^{d}$ with elements $b_{i}^{(i)}=3/2$ and $b_{j}^{(i)}=1/2$ for $j\neq i$, and $R_{-a}(\mathbf{b};\mathbf{z})$ is the Carlson-R function
$R_{-a}(\mathbf{b};\mathbf{z})$ $=\frac{1}{B(a,\sum{b_{i}})}\int_{0}^{\infty}t^{\sum{b_{i}}-a}\prod(t+z_{i})^{-b_{i}}dt$
$=\frac{1}{B(a,\sum{b_{i}})}\int_{0}^{\infty}t^{a-1}\prod(1+tz_{i})^{-b_{i}}dt.$
We use some properties of the Carlson-R function to simplify (a) further
$-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\sum_{i=1}^{d}\frac{1}{d}B(\frac{1}{2},\frac{d+1}{2})S_{ii}R_{-\frac{1}{2}}\left(\mathbf{b}^{(i)};\mathrm{diag}\left(\frac{2}{d}\mathbf{S}\right)\right)$ (54)
$\overset{(b)}{=}$ $-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\sum_{i=1}^{d}\frac{1}{\sqrt{2d}}B(\frac{1}{2},\frac{d+1}{2})S_{ii}R_{-\frac{1}{2}}\left(\mathbf{b}^{(i)};\mathrm{diag}\left(\mathbf{S}\right)\right)$
$\overset{(c)}{=}$ $-\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\sqrt{2d}B(\frac{1}{2},\frac{d+1}{2})\frac{1}{2}R_{\frac{1}{2}}\left(\frac{1}{2}\mathbf{1};\mathrm{diag}\left(\mathbf{S}\right)\right)$
$=$ $-\tilde{\sigma}_{x}\sqrt{2d}\frac{\Gamma({\frac{d+1}{2})}}{\Gamma({\frac{d}{2}+1)}}R_{\frac{1}{2}}\left(\frac{1}{2}\mathbf{1};\mathrm{diag}\left(\mathbf{S}\right)\right);$
the step (b) follows [18, (5.9-3)] that
$R_{-\frac{1}{2}}\left(\mathbf{b}^{(i)};\mathrm{diag}\left(\frac{2}{d}\mathbf{S}\right)\right)=(\frac{2}{d})^{-1/2}R_{-\frac{1}{2}}\left(\mathbf{b}^{(i)};\mathrm{diag}\left(\mathbf{S}\right)\right),$
and the step (c) comes from Euler’s homogeneity relation [18, (5.9-2)]
$\sum^{d}_{i=1}S_{ii}\frac{\partial}{\partial S_{ii}}R_{\frac{1}{2}}\left(\frac{1}{2}\mathbf{1};\mathrm{diag}\left(\mathbf{S}\right)\right)=\frac{1}{2}R_{\frac{1}{2}}\left(\frac{1}{2}\mathbf{1};\mathrm{diag}\left(\mathbf{S}\right)\right),$
and the derivative of R function [18, (5.9-9)]
$\frac{\partial}{\partial S_{ii}}R_{\frac{1}{2}}\left(\frac{1}{2}\mathbf{1};\mathrm{diag}\left(\mathbf{S}\right)\right)=\frac{1}{2d}R_{-\frac{1}{2}}\left(\mathbf{b}^{(i)};\mathrm{diag}\left(\mathbf{S}\right)\right).$
The last equality of (54) comes from
$B(\frac{1}{2},\frac{d+1}{2})=\Gamma(\frac{1}{2})\Gamma(\frac{d+1}{2})/\Gamma(\frac{d}{2}+1),$
and thus, (53) becomes the objective of (27).
Next, with aids of Dirichlet average [18], we prove that the objective of (27) is Schur-convex [19], and it achieves the global minimum when all the diagnol elements of $\mathbf{S}$ are the same as $s$.
Lemma 6.
The objective of equivalent problem (27) of (24) is symmetric and convex over $\mathbf{S}$, with such a Schur-convex objective (27) has global minimum when $\mathbf{S}=s\mathbf{I}$.
Proof.
Carlson R function is a special case of Dirichlet average[18, (5.9-1)]
$R_{\frac{1}{2}}(\frac{1}{2}\mathbf{1};\mathrm{diag}(\mathbf{S}))=\int_{\mathbb{R}^{d-1}}(\mathbf{u}\cdot\mathrm{diag}(\mathbf{S}))^{1/2}d\mu_{\frac{1}{2}\mathbf{1}}(\mathbf{u}),$ (55)
where
$d\mu_{\frac{1}{2}\mathbf{1}}(\mathbf{u}):=\frac{\Gamma(d/2)}{\Gamma(1/2)^{d}}u_{1}^{-1/2}u_{2}^{-1/2}\cdots u_{d}^{-1/2}du_{1}\cdots du_{d-1},$
and
$\sum_{i=1}^{d}u_{i}=1,\ u_{i}\geq 0.$
Note that (55) is symmetric over $\mathbf{S}$ obviously, and thus same property applies to (27). Since the Hessian matrix of $(\mathbf{u}\cdot\mathrm{diag}(\mathbf{S}))^{1/2}$ is positive semidefinite and the nonnegative weighted integral preserves the convexity, (55) is concave and thus (27) is convex over $\mathbf{S}$. ∎
Simplified with Lemma 6, (27) becomes
$\min_{s}s-\tilde{\sigma}_{x}\sqrt{2d}\frac{\Gamma({\frac{d+1}{2})}}{\Gamma({\frac{d}{2}+1)}}R_{\frac{1}{2}}(\frac{1}{2}\mathbf{1};s\cdot\mathbf{1})$ (56)
$=$ $\min_{s}s-\tilde{\sigma}_{x}\sqrt{2d}\frac{\Gamma({\frac{d+1}{2})}}{\Gamma({\frac{d}{2}+1)}}s^{1/2},$
where the equality follows [27, 19.16.11]. Then, by solving the quadratic programming the optimal
$s=\frac{d}{2}\frac{\Gamma(\frac{d+1}{2})^{2}}{\Gamma(\frac{d}{2}+1)^{2}}\tilde{\sigma}_{x}^{2}$
with minimum value of (22) when $q=2$ is
$\tilde{\sigma}_{x}\sqrt{1-\frac{d}{2}\frac{\Gamma(\frac{d+1}{2})^{2}}{\Gamma(\frac{d}{2}+1)^{2}}}.$ (57)
By Lemma 1, the upper bound (26) for (II) from (III-B) is obtained. From [27, (5.11.12)], we have the asymptotic equivalence (same notation as $X\sim\mathcal{N}$)
$\frac{d}{2}\left(\Gamma(\frac{d+1}{2})/\Gamma(\frac{d}{2}+1)\right)^{2}\sim 1.$
The upper bound (26) then goes to zero as $d\rightarrow\infty$, and the lower bound of (II) is zero obviously. Thus, the linear generator with the parameters (19) is asymptotically optimal for the sliced WGAN (II).
IV-C Proof of Lemma 2
The R.H.S of (28) can be solved by following the proof steps in Sec. IV-A. When $q=2$, for any $\omega$,
$\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]$ (58)
$=$ $\int_{\mathbb{R}}|\omega^{T}x-t^{\Theta}_{\omega}(\omega^{T}x)|^{2}d\mu_{\omega}(\omega^{T}x),$
where the optimal trasport function
$t^{\Theta}_{\omega}(\omega^{T}x)=\sup\left\{y\in\mathbb{R}:F_{\nu^{\Theta}_{w}}(y)\leq F_{\mu_{\omega}}\left(\omega^{T}x\right)\right\}.$ (59)
For the L.H.S of (28), when $q=2$, for any $\omega$,
$\inf_{\pi\in\Pi(\mu,\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]$ (60)
$=$ $\int_{\mathbb{R}^{d}}|\omega^{T}x-t^{\Theta}(x)|^{2}d\mu(x),$ (61)
where the optimal transport function $t^{\Theta}(x)$ equals to $t^{\Theta}_{\omega}(\omega^{T}x)$ in (59), which will be proved in Lemma 8 later. Then, with the change of variables[28], (61) and (58) are equal.
To show that the optimal transport function $t^{\Theta}(x)$ for (60) equals to $t^{\Theta}_{\omega}(\omega^{T}x)$ in (59), we first prove that (60) belongs to the general multi-to-one optimal transport problem in [21] by checking its properties. To do this, let $Y=\omega^{T}G_{\Theta}(Z)$ and introduce the following sets. For each fixed $y\in\mathbb{R}$, for all $k\in\mathbb{R}$, we define sets
$X(y,k)$ $:=\left\{x\in\mathbb{R}^{d}:-\frac{\partial}{\partial y}c(x,y)=k\right\}$
$=\left\{x\in\mathbb{R}^{d}:\omega^{T}x=y+{k}/{2}\right\},$
$X_{\leq}(y,k)$ $:=\left\{x\in\mathbb{R}^{d}:-\frac{\partial}{\partial y}c(x,y)\leq k\right\}$
$=\left\{x\in\mathbb{R}^{d}:\omega^{T}x\leq y+{k}/{2}\right\},$ (62)
where $c(x,y)=(\omega^{T}x-y)^{2}$.
Lemma 7.
Suppose that $\mu$ vanishes on hypersurfaces in $\mathbb{R}^{d}$ and $\nu_{w}^{\Theta}$ has no atom. For multi-to-one-dimensional inner discriminator problem (60), we have the following properties
  • (a)
    For each $y\in\mathbb{R}$, there exists a maximal interval $K(y)=[k^{-}(y),k^{+}(y)]\neq\emptyset$ such that
    $\mu(X_{\leq}(y,k))=\nu_{w}^{\Theta}((-\infty,y))\quad\forall k\in K(y).$ (63)
  • (b)
    $\pi=(id\times t^{\Theta}(x))_{\#}\mu$ is the unique minimizer of (60), where $t^{\Theta}(x):\mathbb{R}^{d}\to\mathbb{R}$ is defined by
    $t^{\Theta}(x)=y\quad\hbox{if}\quad x\in X(y,k^{+}(y)).$ (64)
Proof.
(a) $c(x,y)$ is non-degenerate in the sense of [21, Definition 3.1] since the rank of $D^{2}_{xy}c(x,y)\geq 1$ for all $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$. The desired result follows from [21, Theorem 4(a)].
$\frac{\partial^{2}}{\partial x\partial y}c^{w}(x,y)=\begin{bmatrix}2ww^{T}&-2w\\ -2w^{T}&2\end{bmatrix}$.
(b) To apply the optimal transport in [21], we need to check whether our (62) meets the requirements listed within. We first show that $f^{\pm}(y):=y\mapsto y+{k^{\pm}(y)}/{2}$ are non-decreasing. By contradiction, suppose that there exist $y<y^{\prime}$ such that $f^{+}(y)>f^{+}(y^{\prime})$. Observe from (62) that
$X_{\leq}(y,k^{+}(y))=\{x\in\mathbb{R}^{d}:w^{T}x\leq f^{+}(y)\}$ (65)
$\supseteq$ $\{x\in\mathbb{R}^{d}:w^{T}x\leq f^{+}(y^{\prime})\}=X_{\leq}(y^{\prime},k^{+}(y^{\prime})),$
which implies $\mu(X_{\leq}(y,k^{+}(y)))\geq\mu(X_{\leq}(y^{\prime},k^{+}(y^{\prime})))$. On the other hand, (63) entails
$\mu(X_{\leq}(y,k^{+}(y)))=\nu^{\Theta}_{w}((-\infty,y))$ (66)
$\leq$ $\nu^{\Theta}_{w}((-\infty,y^{\prime}))=\mu(X_{\leq}(y^{\prime},k^{+}(y^{\prime}))).$
We therefore conclude
$\mu(X_{\leq}(y,k^{+}(y)))=\mu(X_{\leq}(y^{\prime},k^{+}(y^{\prime})))$
and consequently
$\nu^{\Theta}_{w}((-\infty,y))=\nu^{\Theta}_{w}((-\infty,y^{\prime})).$ (67)
As $y+{k^{+}(y)}/{2}=f^{+}(y)>f^{+}(y^{\prime})=y^{\prime}+{k^{+}(y^{\prime})}/{2}$, we can choose $\bar{k}>k^{+}(y^{\prime})$ such that $y^{\prime}+\bar{k}/2=y+{k^{+}(y)}/{2}$. By (62), this implies $X_{\leq}(y^{\prime},\bar{k})=X_{\leq}(y,k^{+}(y))$. It then follows that
$\mu(X_{\leq}(y^{\prime},\bar{k}))=\mu(X_{\leq}(y,k^{+}(y)))$
$=\nu^{\Theta}_{w}((-\infty,y))=\nu^{\Theta}_{w}((-\infty,y^{\prime})),$
where the second equality follows from (63) and the third equality is due to (67). This readily contradicts part (a), which asserts that $k^{+}(y^{\prime})$ is the largest solution $k\in\mathbb{R}$ to $\mu(X_{\leq}(y^{\prime},k))=\nu^{\Theta}_{w}((-\infty,y^{\prime}))$. A similar argument shows that $y\mapsto f^{-}(y)$ is also non-decreasing. As $y\mapsto y+{k^{\pm}(y)}/{2}$ are non-decreasing, we see from (62) that
$\hbox{$y\mapsto X_{\leq}(y,k^{\pm}(y))$ are non-decreasing}.$ (68)
Now, consider $y<y^{\prime}$ such that $\nu^{\Theta}_{w}((y,y^{\prime}))>0$. With $\nu^{\Theta}_{w}((-\infty,y))<\nu^{\Theta}_{w}((-\infty,y^{\prime}))$, we deduce from (63) that $\mu(X_{\leq}(y,k^{+}(y)))=\nu^{\Theta}_{w}((-\infty,y))<\nu^{\Theta}_{w}((-\infty,y^{\prime}))=\mu(X_{\leq}(y^{\prime},k^{-}(y^{\prime})))$. This, together with (62), indicates that $y+k^{+}(y)/2<y^{\prime}+k^{-}(y^{\prime})/2$. This in turn implies that
$X_{\leq}(y,k^{+}(y))\subseteq X_{<}(y^{\prime},k^{-}(y^{\prime})).$ (69)
Then, the desired result follows from [21, Theorem 4(b)]. ∎
From Lemma 7(b), the optimal transport function in (61) has the closed-form solution (64). We prove that this optimal transport function is well-defined by properties in Lemma 7.
Lemma 8.
For $\mu$ almost every ($\mu-a.e.)\ x\in\mathbb{R}^{d}$, $t^{\Theta}(x)=F_{\nu^{\Theta}_{\omega}}^{-1}(F_{\mu_{\omega}}(\omega^{T}x))$ and can be equivalently expressed as the R.H.S of (59).
Proof.
Note that $F_{\nu^{\Theta}_{\omega}}$ is strictly increasing on
$\mathbb{R}^{\nu^{\Theta}_{\omega}}:=\mathbb{R}\backslash\bigcup_{y<y^{\prime},\ \nu^{\Theta}_{\omega}((y,y^{\prime}))=0}[y,y^{\prime}).$
Hence, $F_{\nu^{\Theta}_{\omega}}^{-1}$ is well-defined on $\mathbb{R}^{\nu^{\Theta}_{\omega}}$. For any $y<y^{\prime}$ such that $\nu^{\Theta}_{\omega}((y,y^{\prime}))=0$, we have $\nu^{\Theta}_{\omega}((-\infty,y))=\nu^{\Theta}_{\omega}((-\infty,y^{\prime}))$. By Lemma 7(a), this implies
$\mu\big{(}\{x\in\mathbb{R}^{d}:y+\frac{k^{+}(y)}{2}<\omega^{T}x\leq y^{\prime}+\frac{k^{+}(y^{\prime})}{2}\}\big{)}=0.$
Since $\mu$ vanishes on hypersurfaces, $\mu\big{(}\{x\in\mathbb{R}^{d}:\omega^{T}x=y+\frac{k^{+}(y)}{2}\})$ for $(y,y^{\prime})$ such that $\nu^{\Theta}_{w}((y,y^{\prime}))=0$ can be ignored. It follows that
$\mu\big{(}\{x\in\mathbb{R}^{d}:\omega^{T}x=y+\frac{k^{+}(y)}{2}\text{for some $y\in\mathbb{R}^{\nu^{\Theta}_{\omega}}$}\}\big{)}=1.$
That is, for $\mu-a.e.\ x\in\mathbb{R}^{d}$, $t^{\Theta}(x)\in\mathbb{R}^{\nu^{\Theta}_{\omega}}$. Since we have the following relationship between CDFs by (62), (63) , and (64),
$F_{\mu_{\omega}}\left(\omega^{T}x\right)=F_{\mu_{\omega}}\left(y+\frac{k^{+}(y)}{2}\right)=F_{\nu^{\Theta}_{w}}(y)=F_{\nu^{\Theta}_{w}}\left(t^{\Theta}(x)\right),$ (70)
then $t^{\Theta}(x)=F_{\nu^{\Theta}_{\omega}}^{-1}(F_{\mu_{\omega}}(\omega^{T}x))$. ∎
V Empirical study on convergence with synthetic data
We have only considered population WGAN thus far. In practice, the distribution is estimated from training data and does not have a closed-form CDF. Therefore, it is of interest to empirically study how fast the solution converges to the population WGAN result using synthetic training data. In Fig. 1, we consider the linear activation and plot the optimal $\theta_{2}^{*}$ in (16)(17) obtained by solving population WGAN with $d=1,q=2$, and its estimate with synthetic data when $\mu$ is chosen to be (a) $\mathcal{N}(0,1)$ and (b) Laplace distribution [29] with mean 0 and scale $1/\sqrt{2}$ (which also has a unit variance). Specifically, by generating training data $\{x_{i}\}^{M}_{i=1}$ according to the true distribution $\mu$, we empirically estimate $\theta_{2}^{*}$ from (16) by
$\frac{1}{M}\sum^{M}_{i=1}x_{i}\Phi^{-1}(\hat{F}_{\mu}(x_{i}))$ (71)
and the kernel density estimation (KDE) [30] is used to replace the true CDF $F_{\mu}(x)$ with the estimated one $\hat{F}_{\mu}(x)$ from the training data. For each distribution, our result shows a nice convergence behavior where the difference becomes smaller than 0.005 with only MM=50000 data size. The optimal $\theta_{2}^{*}$ in (16)(17) is 1 and 0.98013 for Gaussian and Laplace distributions respectively. Using linear generator, GAN output is also Gaussian and thus the convergence behavior for Gaussian data is better than that for Laplace data. Note that $\theta^{*}_{1}$ in (16) can be estimated by the sample mean, so the convergence is not shown in Fig. 1. We also use SGD with momentum to estimate $\theta_{1}$ and $\theta_{2}$ for loss function (1) under $d=1,q=1$, with gradient empirically obtained from (18) using KDE $\hat{F}_{\mu}(x)$. With $\mathcal{N}(\mu=1.5,\sigma^{2}=4)$ data, $(\theta_{1},\theta_{2})$ converges to $(1.4957,2.0905)$ shown in Fig. 2 and close to the population local optimum $(\mu,\sigma)$ for $q=1$.
The convergence rate of the empirical WGAN solution to that of the population WGAN problem has been theoretically analyzed to be $M^{-2/d}$ for the LQG setting in [8, Theorem 2]. Note that the convergence of iteratively solving empirical WGAN based on a good solver for (2) was given in [12]. However, this convergence heavily relies on a good semi-discrete optimal transport solver, which is still hard to design [31]. It is our future work to analyze the convergence rate for our non-Gaussian results, also compare it with the empirical simulations for $d>1$.
Figure 1: Comparison of optimal parameters in (16)(17) with their estimates (71) using synthetic data.
Figure 2: the estimation of $\theta_{1}$ and $\theta_{2}$ in W1Gaussain via SGD
Figure 3: Comparison of sliced Wasserstein distance with our optimal parameter and [8]’s when data is i.i.d.
Figure 4: Comparison of sliced Wasserstein distance with our optimal parameter and [8]’s when data is correlated.
For the sliced WGAN problem, we compare our asymptotically optimal linear generator in theorem 3 with the $r$-PCA solution which is the optimal WGAN solution in the LQG setting. First, We use synthetic data $\{X_{1:d}^{(i)}\}_{i=1}^{M}$ with data size $M=4\times 10^{5}$, where $\{X_{1:d}^{(i)}\}$ generated follows i.i.d Laplace distribution with zero mean and unit variance for all $i$, and estimate the empirical distribution $\mu=M^{-1}\sum_{i=1}^{M}\delta_{X^{(i)}}$. We compute the empirical sliced Wasserstein distance with Monte Carlo based on $2\times 10^{4}$ random projections,
$\left(\sum_{\omega_{i}\in\Omega,\ i=1}^{20000}\inf_{\pi\in\Pi(\mu_{\omega},\nu^{\Theta}_{\omega})}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|^{2}]d\omega\right)^{\frac{1}{2}}.$ (72)
Then, we let $\Theta$ of (72) be our proposed linear generator parameters (19) and the parameters of $r$-PCA solution. In Fig. 3, we can see that the sliced Wasserstein distance with our proposed linear generator parameters is much lower than the one of $r$-PCA when $r=\frac{2}{3}d$.
In addition, we use not only the i.i.d. synthetic data but also the data generated by AR model mentioned in Remark 1. The AR model is defined as $X_{j}=0.5X_{j-1}+\eta_{j}$ where $\eta_{j}$ is i.i.d. student distribution noise. Given the correlated data, we compare the value of (72) with our proposed linear generator parameters to the value with the parameters of $r$-PCA. In Fig. 4, except for $d=10$, we can see that the sliced Wasserstein distance with our proposed linear generator parameters is lower like the i.i.d. data scenario when $r=\frac{2}{3}d$.
When $r=d$, our results reach the same value as $r$-PCA method with both i.i.d. data and correlated data at $d=1000$ and $d=2000$, respectively. Although $r$-PCA method can have lower sliced Wasserstein distance in some case like $r=d$, the computation complexity of solving the $r$-PCA optimizer is higher than the complexity of our asympotically optimizer. For the $r$-PCA optimizer, the complexity of computing the covariance and find the eigenvalues is $O((M+r)d^{2})$. For our proposed linear generator parameters, we only need to compute $\mathbb{E}_{\mu}[\|X\|^{2}]$ which complexity is $O(Md)$. Our asymptotically optimal linear generator has comparable performance to the r-PCA solution while reducing the computation energy resources.
Acknowledgement
We would like to acknowledge the help of Wen-Yi Tseng and Kuan-Hui Lyu on the simulations.
Appendix A Proof for Theorem 2
As $\Psi$ is neither continuous nor strictly increasing, (31) does not hold in general and we cannot directly apply proofs in Sec. IV-A. Again assume $\theta_{2}\in\mathbb{R}_{+}$, Under the condition $\theta_{2}\geq 0$, we deduce from (29) and (12) that
$t^{\Theta}(x)=\begin{cases}\theta_{1},\quad&\hbox{if}\ F_{\mu}(x)\leq 1/2,\\ \theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x)),\quad&\hbox{if}\ F_{\mu}(x)>1/2.\end{cases}$ (73)
Similarly to (35), the sub-problem (33) now becomes
$\begin{split}&\min_{\theta_{1},\theta_{2}\in\mathbb{R}}J(\theta_{1},\theta_{2}):=\int_{\{F_{\mu}(x)\leq 1/2\}}\Big{(}\theta_{1}-x\Big{)}^{2}F_{\mu}^{\prime}(x)dx\\ &\hskip 21.68121pt+\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big{)}^{2}F_{\mu}^{\prime}(x)dx\\ &\hbox{subject to}\quad g(\theta_{1},\theta_{2}):=-\theta_{2}\leq 0.\end{split}$ (74)
The corresponding first-order condition—namely, the KKT condition—is again in the form of (36)-(37), where $\lambda\geq 0$ is the Lagrange multiplier. By direct calculation, (36) becomes
$\int_{\{F_{\mu}(x)\leq 1/2\}}\Big{(}\theta_{1}-x\Big{)}F_{\mu}^{\prime}(x)dx$ (75)
$+$ $\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big{)}F_{\mu}^{\prime}(x)dx=0,$
$2\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(x))-x\Big{)}$ (76)
$\Phi^{-1}(F_{\mu}(x))F_{\mu}^{\prime}(x)dx-\lambda=0.$
Let $X$ be a random variable with cdf $F_{\mu}$. Then, (75)-(76) can be written as
$\theta_{1}+\theta_{2}\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]-\mathbb{E}[X]=0,$ (77)
$\theta_{1}\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]$ (78)
$+$ $\theta_{2}\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)>1/2\}}\right]$
- $\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]-\lambda/2=0.$
Plugging $\theta_{1}=\mathbb{E}[X]-\theta_{2}\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]$ (from (77)) into (78), we obtain
$\mathbb{E}[X]\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]$
$+$ $\theta_{2}\text{Var}\left(\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)$
- $\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]-\lambda/2=0.$ (79)
Hence, a solution $(\theta_{1},\theta_{2},\lambda)$ to (36) must equivalently satisfy (77) and (79). Note that $F_{\mu}(X)\sim\operatorname{Uniform}[0,1]$, so that the cdf of $\Phi^{-1}(F_{\mu}(X))$ is simply $\Phi$. That is, $\Phi^{-1}(F_{\mu}(X))$ and $Z$ have identical distribution. It follows that
$\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right]$
$=$ $\mathbb{E}[Z1_{\{\Phi(Z)>1/2\}}]=\mathbb{E}[Z1_{\{Z>0\}}]=\frac{1}{\sqrt{2\pi}},$ (80)
$\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)>1/2\}}\right]$
$=$ $\mathbb{E}[Z^{2}1_{\{\Phi(Z)>1/2\}}]=\mathbb{E}[Z^{2}1_{\{Z>0\}}]=\frac{1}{2}.$ (81)
In view of this, (77) and (79) reduce to
$\begin{split}\theta_{1}&=\mathbb{E}[X]-\frac{\theta_{2}}{\sqrt{2\pi}},\\ \lambda/2&=\frac{\theta_{2}}{2}\left(1-\frac{1}{\pi}\right)-\text{Cov}(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}).\end{split}$ (82)
Lemma 9.
Suppose that $\mu$ has no atom. Then, (74) has a unique minimizer $(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$ given by
$\begin{split}\theta^{*}_{1}&=\mathbb{E}[X]-\frac{\sqrt{2\pi}}{\pi-1}\operatorname{Cov}\big{(}X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\big{)},\\ \theta^{*}_{2}&=\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)\geq 0.\end{split}$ (83)
Proof.
As $\Phi^{-1}$ is strictly increasing and $F_{\mu}$ is nondecreasing, the map $x\mapsto\Phi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}$ is nondecreasing. Hence, by a similar argument in (40), $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)\geq 0$. The result then follows by repeating the same arguments in the proof of Lemma 1. Specifically, we deal with the two cases where $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)$ is strictly positive and zero separately. In each case, we solve (82) (in place of (38) in Lemma 1) and find that $H_{J}+\lambda H_{g}$ now takes the form
$\begin{bmatrix}1&\frac{1}{\sqrt{2\pi}}\\ \frac{1}{\sqrt{2\pi}}&\frac{1}{2}\end{bmatrix}$,
which is positive definite. ∎
Lemma 10.
Suppose that $\mu$ has no atom and $h(x)=\max\{x,0\}$ for all $x\in\mathbb{R}$. The sub-problem (33) has a unique minimizer $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$, defined as in (83). Moreover, the minimum value is
$\begin{split}&\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)\\ =&\operatorname{Var}(X)-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)>1/2\}}\right)^{2}.\end{split}$ (84)
Proof.
As (33) is equivalent to (74), Lemma 9 directly shows that $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})\in\mathbb{R}\times\mathbb{R}_{+}$ as in (83) is the unique minimizer of (33). Thanks to (73) and (83),
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$
$=$ $\int_{\{F_{\mu}(x)\leq 1/2\}}\left(x-\mathbb{E}[X]+\frac{\theta^{*}_{2}}{\sqrt{2\pi}}\right)^{2}d\mu(x)$
$+$ $\int_{\{F_{\mu}(x)>1/2\}}\left(x-\mathbb{E}[X]-\theta^{*}_{2}(\Psi^{-1}(F_{\mu}(x))-\frac{1}{\sqrt{2\pi}})\right)^{2}d\mu(x)$
$=$ $\int_{\{F_{\mu}(x)\leq 1/2\}}(x-\mathbb{E}[X])^{2}d\mu(x)$
$+$ $\sqrt{\frac{2}{\pi}}\theta^{*}_{2}\int_{\{F_{\mu}(x)\leq 1/2\}}(x-\mathbb{E}[X])d\mu(x)$
$+$ $\frac{(\theta^{*}_{2})^{2}}{2\pi}\int_{\mathbb{R}}1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$+$ $\int_{\{F_{\mu}(x)>1/2\}}(x-\mathbb{E}[X])^{2}d\mu(x)$
- $2\theta^{*}_{2}\int_{\{F_{\mu}(x)>1/2\}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))-\frac{1}{\sqrt{2\pi}}\Big{)}d\mu(x)$
$+$ $(\theta^{*}_{2})^{2}\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\Psi^{-1}(F_{\mu}(x))-\frac{1}{\sqrt{2\pi}}\Big{)}^{2}d\mu(x).$ (85)
Observe that
$\int_{\{F_{\mu}(x)>1/2\}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))-\frac{1}{\sqrt{2\pi}}\Big{)}d\mu(x)$
$=$ $\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}$
- $\frac{1}{\sqrt{2\pi}}1_{\{F_{\mu}(x)>1/2\}}\Big{)}d\mu(x)$
$=$ $\int_{\mathbb{R}}(x-\mathbb{E}[X])\Big{(}\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\frac{1}{\sqrt{2\pi}}\Big{)}d\mu(x)$
$+$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}(x-\mathbb{E}[X])1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$=$ $\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}\right)$
$+$ $\frac{1}{\sqrt{2\pi}}\int_{\mathbb{R}}(x-\mathbb{E}[X])1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x),$ (86)
where the last equality follows from (80). Similarly,
$\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\Psi^{-1}(F_{\mu}(x))-\frac{1}{\sqrt{2\pi}}\Big{)}^{2}d\mu(x)$
$=$ $\int_{\mathbb{R}}\Big{(}\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\frac{1}{\sqrt{2\pi}}1_{\{F_{\mu}(x)>1/2\}}\Big{)}^{2}d\mu(x)$
$=$ $\int_{\mathbb{R}}\Big{(}\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}-\frac{1}{\sqrt{2\pi}}\Big{)}^{2}d\mu(x)$
- $\frac{1}{2\pi}\int_{\mathbb{R}}1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$=$ $\operatorname{Var}\left(\Psi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)>1/2\}}\right)-\frac{1}{2\pi}\int_{\mathbb{R}}1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x)$
$=$ $\frac{\pi-1}{2\pi}-\frac{1}{2\pi}\int_{\mathbb{R}}1_{\{F_{\mu}(x)\leq 1/2\}}d\mu(x),$ (87)
where the last two equalities follow from (80) and (81). Thanks to (86) and (87), we conclude from (85) that
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$
$=$ $\operatorname{Var}(X)-2\theta^{*}_{2}\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(x)>1/2\}}\right)$
$+(\theta^{*}_{2})^{2}\frac{\pi-1}{2\pi}$
$=$ $\operatorname{Var}(X)-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Psi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(x)>1/2\}}\right)^{2},$
where the last equality is due to the definition of $\theta^{*}_{2}$ in (83). ∎
Next, we goes to $\theta_{2}\leq 0$. Under the condition $\theta_{2}\leq 0$, we deduce from (29) and (12) that
$t^{\Theta}(x)=\begin{cases}\theta_{1}-\theta_{2}\Phi^{-1}(F_{\mu}(x)),\quad&\hbox{if}\ F_{\mu}(x)\leq 1/2,\\ \theta_{1},\quad&\hbox{if}\ F_{\mu}(x)>1/2.\end{cases}$ (88)
Indeed, for the case $F_{\mu}(x)\leq 1/2$, by a calculation similar to (32), but with $\theta_{2}\leq 0$, we get $\Psi\big{(}\frac{t^{\Theta}(x)-\theta_{1}}{\theta_{2}}\big{)}=1-F_{\mu}(x)$, which implies
$\begin{split}&t^{\Theta}(x)=\theta_{1}+\theta_{2}\Psi^{-1}(1-F_{\mu}(x))\\ =&\theta_{1}+\theta_{2}\Phi^{-1}(1-F_{\mu}(x))=\theta_{1}-\theta_{2}\Phi^{-1}(F_{\mu}(x)),\end{split}$
where we used the fact $\Psi(v)=\Phi(v)$ for $v\geq 0$ (recall (12)) and that $Z$ (whose cdf is $\Phi$) is distributed symmetrically about 0.
In the following, we will often write $\theta_{2}\leq 0$ in terms of $\eta:=-\theta_{2}\geq 0$, so that we can more easily borrow the arguments from $\theta_{2}\geq 0$. Similarly to (74), the sub-problem (44) now becomes
$\begin{split}&\min_{\theta_{1}\in\mathbb{R},\eta\geq 0}J(\theta_{1},\eta):=\int_{\{F_{\mu}(x)>1/2\}}\Big{(}\theta_{1}-x\Big{)}^{2}F_{\mu}^{\prime}(x)dx\\ &+\int_{\{F_{\mu}(x)\leq 1/2\}}\Big{(}\theta_{1}+\eta\Phi^{-1}(F_{\mu}(x))-x\Big{)}^{2}F_{\mu}^{\prime}(x)dx\\ &\hbox{subject to}\quad g(\theta_{1},\eta):=-\eta\leq 0.\end{split}$ (89)
The corresponding first-order condition—namely, the KKT condition—is again in the form of (36)-(37) (with $\theta_{2}$ therein replaced now by $\eta$), where $\lambda\geq 0$ is the Lagrange multiplier. Now, by following the same arguments in (75)-(78) (with $\theta_{2}$ replaced by $\eta$ and the two sets $\{F_{\mu}(x)\leq 1/2\}$ and $\{F_{\mu}(x)>1/2\}$ switched), we obtain $\theta_{1}=\mathbb{E}[X]-\eta\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]$ and
$\mathbb{E}[X]\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]$
$+$ $\eta\text{Var}\left(\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)$
- $\mathbb{E}\left[X\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]-\lambda/2=0.$ (90)
Similarly to (80) and (81), we have
$\mathbb{E}\left[\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right]$
$=$ $\mathbb{E}[Z1_{\{\Phi(Z)\leq 1/2\}}]=\mathbb{E}[Z1_{\{Z\leq 0\}}]=-\frac{1}{\sqrt{2\pi}},$ (91)
$\mathbb{E}\left[(\Phi^{-1}(F_{\mu}(X)))^{2}1_{\{F_{\mu}(X)\leq 1/2\}}\right]$
$=$ $\mathbb{E}[Z^{2}1_{\{\Phi(Z)\leq 1/2\}}]=\mathbb{E}[Z^{2}1_{\{Z\leq 0\}}]=\frac{1}{2}.$ (92)
In view of this, (77) and (90) reduce to
$\begin{split}\theta_{1}&=\mathbb{E}[X]+\frac{\eta}{\sqrt{2\pi}},\\ \lambda/2&=\frac{\eta}{2}\left(1-\frac{1}{\pi}\right)-\text{Cov}(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}).\end{split}$ (93)
Lemma 11.
Suppose that $\mu$ has no atom. Then, (89) has a unique minimizer $(\theta_{1}^{*},\eta^{*})\in\mathbb{R}\times\mathbb{R}_{+}$ given by
$\begin{split}\theta^{*}_{1}&=\mathbb{E}[X]+\frac{\sqrt{2\pi}}{\pi-1}\operatorname{Cov}\big{(}X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\big{)},\\ \eta^{*}&=\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)\geq 0.\end{split}$ (94)
Proof.
Consider $f(x):=\Phi^{-1}(F_{\mu}(x))1_{\{F_{\mu}(x)\leq 1/2\}}$ and set $x^{*}:=\sup\{x\in\mathbb{R}:F_{\mu}(x)\leq 1/2\}$. As $F_{\mu}$ is nondecreasing and $\Phi^{-1}$ is strictly increasing, for $x<x^{*}$, $f(x)=\Phi^{-1}(F_{\mu}(x))$ is nondecreasing with $f(x)\leq\Phi^{-1}(1/2)=0$; for $x\geq x^{*}$, we simply have $f\equiv 0$. Since $f$ is nondecreasing on $\mathbb{R}$, an argument similar to (40) implies $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)\geq 0$. The result then follows by repeating the same arguments in the proof of Le,,a 1. Specifically, we deal with the two cases where $\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)$ is strictly positive and zero separately. In each case, we solve (93) (in place of (38) in Lemma 1) and find that $H_{J}+\lambda H_{g}$ now takes the form
$\begin{bmatrix}1&-\frac{1}{\sqrt{2\pi}}\\ -\frac{1}{\sqrt{2\pi}}&\frac{1}{2}\end{bmatrix}$,
which is positive definite. ∎
Lemma 12.
Suppose that $\mu$ has no atom and $h(x)=\max\{x,0\}$ for all $x\in\mathbb{R}$. The sub-problem (44) has a unique minimizer $\Theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})$, defined by
$\begin{split}\theta^{*}_{1}&=\mathbb{E}[X]+\frac{\sqrt{2\pi}}{\pi-1}\operatorname{Cov}\big{(}X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\big{)},\\ \theta^{*}_{2}&=-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right).\end{split}$ (95)
Moreover, the minimum value is
$\int_{\mathbb{R}}|x-t^{\Theta^{*}}(x)|^{2}d\mu(x)$ (96)
$=$ $\operatorname{Var}(X)-\frac{2\pi}{\pi-1}\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))1_{\{F_{\mu}(X)\leq 1/2\}}\right)^{2}.$
Proof.
Since (44) is equivalent to (89) by taking $\theta_{2}=-\eta$, Lemma 11 directly provides the unique minimizer $\theta^{*}=(\theta_{1}^{*},\theta_{2}^{*})$, given as in (95), for (44). By repeating all the calculations in Lemma 10 (with the two sets $\{F_{\mu}(x)\leq 1/2\}$ and $\{F_{\mu}(x)>1/2\}$ switched and $\theta_{2}$ replaced by $\eta^{*}$ in (94)), we obtain the minimum value formula. ∎
Finally we can get our theorem. If (13) holds, the value in (84) is smaller or equal to that in (96). Hence, the minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should be the one associated with the value (84), which is given by (83). On the other hand, if (13) fails, the value in (96) is smaller than that in (84). The minimizer $(\theta_{1}^{*},\theta_{2}^{*})$ should then be the one associated with the value (96), which is given by (95).
Appendix B Proof of Corollary 1
Besides checking (8), it is easy to see that one can limit $\theta_{2}\in\mathbb{R}_{+}$ such that the optimal parameter is (9) when $h(Z)=Z$. Now the distribution of h(Z)-h(Z) is still the same as $h(Z)$. Then one can rewrite (7) as $G_{\Theta}(Z)=\theta_{1}+(-\theta_{2})*(-Z),$ and absorb the case for $\theta_{2}<0$ into that for $\theta_{2}\geq 0$. Since now $h(Z)=Z,Z\sim\mathcal{N}(0,1)$, (9) reduces to
$\begin{split}\theta^{*}_{2}&=\operatorname{Cov}\left(X,\Phi^{-1}(F_{\mu}(X))\right)\\ \theta^{*}_{1}&=\mathbb{E}_{\mu}[X];\end{split}$ (97)
Furthermore,
$\Phi^{-1}(F_{\mu}(X))\stackrel{{\scriptstyle d}}{{=}}Z,$ (98)
from $F_{\mu}(X)\stackrel{{\scriptstyle d}}{{=}}\Phi(Z)$ and both uniformly distributed in $[0,1]$ for continuous $X$. Thus $\mathbb{E}[\Phi^{-1}(F_{\mu}(X))]=0$ and we reach (16). The other method is that one can rewrite (1) as
$\min_{\Theta}\inf_{\pi\in\Pi(\mu,\nu)}\int\|x-G_{\Theta}(z)\|^{q}d\pi(x,z).$ (99)
where $\nu$ is the distribution of generator input, a Gaussian vector. The reason is by transform of random variables, the optimal solution the inner discriminator problem of (1) can be transformed to be the optimal solution of inner discriminator problem in (99). Let $\pi^{*}(x,y,\Theta)$ be the solution of the inner discriminator problem, then from Danskin’s theorem [32, Proposition B.25], its gradient over $\Theta$ is
$\int\nabla_{\!\Theta}\|G_{\Theta}(z)-x\|^{q}d\pi^{*}(x,z,\Theta),$ (100)
Under W2 with $d=1$, the gradient over vector $\theta=(\theta_{1},\theta_{2})$
$\int\nabla_{\!\Theta}\left(x-(\theta_{1}+\theta_{2}z)\right)^{2}d\pi^{*}(x,z,\theta)$
$=$ $-2\;[\mathbb{E}_{\mu}[X]-(\theta_{1}+\theta_{2}\mathbb{E}[Z])\;,$ (101)
$\mathbb{E}_{\pi*}[XZ]-(\theta_{1}\mathbb{E}[Z]+\theta_{2}\mathbb{E}[Z^{2}])]^{T},$
since from matrix derivative
$\nabla_{\!\Theta}\left(x-(\theta_{1}+\theta_{2}z)\right)^{2}$
[…]
Note that the optimal joint PDF in the LHS of (30) will make $G_{\Theta}(Z)=t^{\Theta}(X)$, which also holds for $\pi^{*}$ in (100). Together with (31)(7), $\pi^{*}$ will make
$\theta_{1}+\theta_{2}\Phi^{-1}(F_{\mu}(X))=\theta_{1}+\theta_{2}Z,$
which results (16) by letting (B) equals to $0,0]^{T}$.
For (17) we swapping the role of $X$ and $G_{\Theta}(Z)$ in Lemma 3, and from (B)
$\theta^{*}_{2}=\mathbb{E}_{\pi*}[XZ]=\mathbb{E}_{g}[F_{\mu}^{-1}(\Phi(Z)\cdot Z].$ (103)
To see this, from Lemma 3
$\inf_{\pi\in\Pi(\mu,\nu^{\Theta})}\int_{\mathbb{R}\times\mathbb{R}}|x-y|^{2}d\pi(x,y)=\int_{\mathbb{R}}|t^{\Theta}(y)-y|^{2}d\nu^{\Theta}(y),$
where $d\nu^{\theta}(y)$ is the distribution of Gaussian variable
$Y=G_{\Theta}(Z)=\theta_{1}+\theta_{2}Z.$ (104)
The optimal transport is then
$t^{\Theta}(y)=F_{\mu}^{-1}\left(\Phi\left(\frac{y-\theta_{1}}{\theta_{2}}\right)\right).$
Note that the optimal joint distribution $\pi*$ will equal to that of $(t^{\Theta}(Y),Y)$. From (104) and $t^{\Theta}(Y)=X$ we reach (103).
Appendix C Detailed Proof for Lemma 1
First we prove that for $q=1,2$, the gap between (21) and (22) is bounded as
$\Bigg{|}$ $\Big{(}\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\Big{)}^{\frac{1}{q}}$-
$\Big{(}\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}G_{\Theta}(Z)|^{q}]d\omega\Big{)}^{\frac{1}{q}}\Bigg{|}\leq f^{\mu}(d)$ (105)
where $f^{\mu}(d)$ is in (48). To see this, first we prove
$\int_{\omega\in\Omega_{G}}W^{q}_{q}(\omega^{T}X,\hat{X}_{G})d\omega\leq(f^{\mu}(d))^{q},q=1,2$
where as the RHS of (2) we define Wasserstein distance of order $q$ as
$W_{q}(\mu,\nu):=\left(\inf_{\pi\in\Pi(\mu,\nu)}\mathbb{E}_{\pi}[\|X-Y\|^{q}]\right)^{1/q}$
where the marginals with respect to the first and second variables are given by $\mu$ and $\nu$ respectively. The case for $q=2$ is proved by [17, Corollary 3]. Now for $q=1$,
$\int_{\omega\in\Omega_{G}}W_{1}(\omega^{T}X,\hat{X}_{G})d\omega\leq\int_{\omega\in\Omega_{G}}W_{2}(\omega^{T}X,\hat{X}_{G})d\omega$
$\leq$ $\bigg{(}\int_{\omega\in\Omega_{G}}W^{2}_{2}(\omega^{T}X,\hat{X}_{G})d\omega\bigg{)}^{1/2}\leq f^{\mu}(d)$
where the first inequality is from $W_{1}\leq W_{2}$ by the definition of Wasserstein distances (see [33, Remark 6.6]) and the second one is from Jensen’s inequality. Then from triangle inequality, given $\omega$
$|W_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))-W_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))|$
$\leq$ $W_{q}(\omega^{T}X,\hat{X}_{G}),q=1,2,$
which implies
$\int_{\omega\in\Omega_{G}}|W_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))-W_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))|^{q}d\omega$
$\leq(f^{\mu}(d))^{q}.$ (106)
Finally, by triangle inequality
$\Bigg{|}$ $\left(\int_{\omega\in\Omega_{G}}W^{q}_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))\right)^{1/q}$
$-\left(\int_{\omega\in\Omega_{G}}W^{q}_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))\right)^{1/q}\Bigg{|}$
$\leq$ $\left(\int_{\omega\in\Omega_{G}}|W_{q}(\omega^{T}X,\omega^{T}G_{\Theta}(Z))-W_{q}(\hat{X}_{G},\omega^{T}G_{\Theta}(Z))|^{q}d\omega\right)^{1/q}.$
Note that when $q=1$, one do not need triangle inequality to make the upper-bound valid. Then with (106), we have (105).
Next, now from $q=1$ in (22) we aim to solve with the linear generator
$\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}\Theta Z|]d\omega\right),$ (107)
which equals to (23) with $q=1$ as follows. Given $\Theta$ and $\omega$, following the proof of Corollary 2
$\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}\Theta Z|]=\frac{1}{\tilde{\sigma}_{x}}\mathbb{E}[|(\tilde{\sigma}_{x}-\sqrt{\omega^{T}\Theta\Theta^{T}\omega})\hat{X}_{G}|],$
since the optimal transport function will make the distribution of $\omega^{T}\Theta Z$ same as
$\frac{\sqrt{\omega^{T}\Theta\Theta^{T}\omega}}{\tilde{\sigma}_{x}}\hat{X}_{G}.$
Then solving (107) equals to solving
$\min_{\Theta}\mathbb{E}_{\omega}[|\tilde{\sigma}_{x}-\sqrt{\omega^{T}\Theta\Theta^{T}\omega}|]$ (108)
since
$\mathbb{E}[|(\tilde{\sigma}_{x}-\sqrt{\omega^{T}\Theta\Theta^{T}\omega})\hat{X}_{G}|]=|\tilde{\sigma}_{x}-\sqrt{\omega^{T}\Theta\Theta^{T}\omega}|\mathbb{E}[|\hat{X}_{G}|].$
Similarly, from $q=2$ in (22) , we aim to solve
$\min_{\Theta}\left(\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\omega^{T}\Theta Z|^{2}]d\omega\right)^{\frac{1}{2}},$ (109)
which equals to solving
$\min_{\Theta}\int_{\omega\in\Omega_{G}}\left(\tilde{\sigma}_{x}-\sqrt{\omega^{T}\Theta\Theta^{T}\omega}\right)^{2}d\omega$ (110)
from Corollary 1. And (110) equals to (23) with $q=2$.
Finally, it is easy to see that (23) with $q=2$ equals to (49). Here we focus on the term
$\mathbb{E}_{w}[\sqrt{\omega^{T}\Theta\Theta^{T}\omega}]=\mathbb{E}_{U_{w}}[U_{w}U_{w}^{-1/2}]$ (111)
in (49) and prove that it equals to (50). By recalling the gamma function $\Gamma(t):=\int_{0}^{\infty}y^{t-1}e^{-y}dy$ for all $t>0$, we observe that
$U_{w}^{-1/2}$ $=U_{w}^{-1/2}\frac{1}{\Gamma(1/2)}\int_{0}^{\infty}y^{1/2-1}e^{-y}dy$
$=\frac{1}{\Gamma(1/2)U_{w}}\int_{0}^{\infty}\bigg{(}\frac{y}{U_{w}}\bigg{)}^{1/2-1}e^{-y}dy$
$=\frac{1}{\Gamma(1/2)}\int_{0}^{\infty}z^{-1/2}e^{-zU_{w}}dz,$
where the second line follows from the change of variable $z=y/U_{w}$. Note that the above calculation is motivated by [34, Example 3.2b.2]. It follows that (111) equals to
$\frac{1}{\Gamma(1/2)}\int^{\infty}_{0}z^{-1/2}\mathbb{E}_{U_{w}}[U_{w}\exp(-zU_{w})]dz$
which equals to (50) from $\mathbb{E}_{U_{w}}[U_{w}\exp(-zU_{w})]=-\frac{\partial}{\partial z}M_{U_{w}}(-z)$ where we recall $M_{U_{w}}(z):=\mathbb{E}[e^{zU_{w}}]$ is the moment generating function of $U_{w}$. Moreover, from [34, Theorem 3.2a.1], $M_{U_{w}}(z)$ has a closed form as
$M_{U_{w}}(z)=\left|\mathbf{I}-\frac{2z}{d}\Theta\Theta^{T}\right|^{-1/2}.$ (112)
Plugging (112) into (50), together with (49)(51), we have (24).
Appendix D Extension of Lemma 1 for $q=1$
To get a solution as (24) for $\Theta$ under $q=1$, we sacrifice the approximation error in Lemma 1 by modifying (105) (106) to get upcoming (116). Details are as follows. Given $\omega$ and $\Theta$, from triangle inequality
$|W_{1}(\omega^{T}X,\omega^{T}\Theta Z)-W_{1}(\hat{X}_{G},\hat{Z}_{G})|$
$\leq$ $W_{1}(\omega^{T}X,\hat{X}_{G})+W_{1}(\omega^{T}\Theta Z,\hat{Z}_{G}),$
where $\hat{Z}_{G}\sim\mathcal{N}(0,\tilde{\sigma}^{2}_{\Theta})$ is independent of $\omega$ with
$\tilde{\sigma}^{2}_{\Theta}=\frac{\mathbb{E}[\|\Theta Z\|^{2}]}{d}=\frac{\mathrm{Tr}(\Theta\Theta^{T})}{d}.$
Then (106) is modified as
$\int_{\omega\in\Omega_{G}}|W_{1}(\omega^{T}X,\omega^{T}\Theta Z)-W_{1}(\hat{X}_{G},\hat{Z}_{G})|d\omega$
$\leq$ $f^{\mu}(d)+f^{\Theta Z}(d),$ (113)
where $f^{\Theta Z}(d)$ is similarly defined as (48) by replacing $\mu$ with $\Theta Z$ and still $O(d^{-1/8})$. Then for every $\Theta$, following previous steps to prove (105) from (106) we have
$\Bigg{|}$ $\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\omega^{T}X-\omega^{T}\Theta Z|]d\omega$ (114)
$-\int_{\omega\in\Omega_{G}}\inf_{\pi}\mathbb{E}_{\pi}[|\hat{X}_{G}-\hat{Z}_{G}|]d\omega\Bigg{|}$
is bounded by $O(d^{-1/8})$ for large $d$. Following the proof of reaching (108), now we aim to solve
$\min_{\Theta}|\tilde{\sigma}_{x}-\sqrt{\mathrm{Tr}(\Theta\Theta^{T})/d}|].$ (115)
and $\Theta$ meeting
$\tilde{\sigma}^{2}_{x}\mathbf{I}=\Theta\Theta^{T}$ (116)
will be the solution.
Appendix E Solution for (22) when $q=d=2$
Corollary 3.
The minimum value of (22) when $q=d=2$ is
$\tilde{\sigma}_{x}\sqrt{1-\frac{\pi}{4}},$
with the optimal parameters of linear generator $\Theta=\mathbf{U}\Lambda\mathbf{V}^{T}$ where $\Lambda\Lambda^{T}=(\pi\tilde{\sigma}_{x}^{2}/4)\mathbf{I},$ and $\mathbf{U},\mathbf{V}^{T}$ are unitary.
Proof.
When $d=2$ the target function (53) equals to
$f(\mathbf{S})=\frac{S_{11}+S_{22}}{2}$-
$\frac{2\tilde{\sigma}_{x}}{\Gamma(1/2)}\int^{\infty}_{0}\frac{1}{2}z^{-1/2}\frac{S_{11}(1+zS_{22})+S_{22}(1+zS_{11})}{[(1+zS_{11})(1+zS_{22})]^{3/2}}dz.$ (117)
We first prove the Schur-convexity of (117) when $d=2$, Recalling the definition of Schur-convexity as follows, our target function $f(\mathbf{S}):\{\mathbb{R}^{+}\cup 0\}^{2}\rightarrow\mathbb{R}$ is Schur-convex if and only if $f$ is symmetric on $\{\mathbb{R}^{+}\cup 0\}^{2}$, and
$(S_{11}-S_{22})\left[\frac{\partial f(\mathbf{S})}{\partial S_{11}}-\frac{\partial f(\mathbf{S})}{\partial S_{22}}\right]\geq 0.$ (118)
To prove (118), we first find the closed-form of $f(\mathbf{S})$ using elliptic integrals as follows. Let us define
$K(m):=\int^{\frac{\pi}{2}}_{0}\frac{d\theta}{\sqrt{1-m\sin^{2}\theta}}$
which is the complete elliptic integral of the first kind, and
$E(m)=\int^{\frac{\pi}{2}}_{0}\sqrt{1-m\sin^{2}\theta}d\theta$
which is the complete elliptic integral of the second kind. When $D_{11}\neq D_{22}\neq 0$, the integral in (117) becomes
$\int_{0}^{\infty}\frac{1}{2}z^{-\frac{1}{2}}\frac{S_{11}(1+zS_{22})}{[(1+zS_{11})(1+zS_{22})]^{\frac{3}{2}}}dz$
$\overset{(a)}{=}$ $\int_{0}^{\infty}\frac{S_{11}(1+x^{2}S_{22})}{[(1+x^{2}S_{11})(1+x^{2}S_{22})]^{\frac{3}{2}}}dx,$
$\overset{(b)}{=}$ $\sqrt{S_{11}}\int_{0}^{\frac{\pi}{2}}\frac{\sec^{2}\theta}{\sec^{3}\theta\sqrt{1+\frac{S_{22}}{S_{11}}\tan^{2}\theta}}d\theta,$
$=$ $\sqrt{S_{11}}\int_{0}^{\frac{\pi}{2}}\frac{\cos^{2}\theta}{\sqrt{\cos^{2}\theta+\frac{S_{22}}{S_{11}}\sin^{2}\theta}}d\theta$
$=$ $\sqrt{S_{11}}\int_{0}^{\frac{\pi}{2}}\frac{1-\sin^{2}\theta}{\sqrt{1+(\frac{S_{22}}{S_{11}}-1)\sin^{2}\theta}}d\theta$
$=$ $\sqrt{S_{11}}\frac{1}{1-\frac{S_{22}}{S_{11}}}\int_{0}^{\frac{\pi}{2}}\frac{-\frac{S_{22}}{S_{11}}+1-(1-\frac{S_{22}}{S_{11}})\sin^{2}\theta}{\sqrt{1-(1-\frac{S_{22}}{S_{11}})\sin^{2}\theta}}d\theta$
$=$ $\frac{-\sqrt{S_{11}}S_{22}}{S_{11}-S_{22}}K(1-\frac{S_{22}}{S_{11}})+\frac{\sqrt{S_{11}}S_{11}}{S_{11}-S_{22}}E(1-\frac{S_{22}}{S_{11}}),$
where change of variable $x=z^{2}$ and $x=\frac{\tan\theta}{\sqrt{D_{11}}}$ are respectively be applied to step (a) and (b). Substituting this results into (117),
[…]
It is easy to check (118) is valid besides the first case in (119). Now we prove that (118) is also valid when $S_{11}\neq S_{22}\neq 0$. Using [27, equation 19.4.1 and equation 19.7.2]
$(S_{11}-S_{22})[\frac{\partial f(S_{11},S_{22})}{\partial S_{11}}-\frac{\partial f(S_{11},S_{22})}{\partial S_{22}}]$ (120)
$=$ $\frac{\tilde{\sigma}_{x}}{\Gamma(1/2)\sqrt{S_{22}}}[(S_{11}+S_{22})K(1-\frac{S_{11}}{S_{22}})-2S_{22}E(1-\frac{S_{11}}{S_{22}})].$
Let
$m=1-\frac{S_{11}}{S_{22}}$
with $m<1$, then proving (120) is non-negative is equivalent to proving
(2m)K(m)2E(m)(2-m)K(m)-2E(m)
is non-negative. For $0\leq m<1$, using [27, equation 19.5.1 and equation 19.5.2],
$(2-m)K(m)-2E(m)$ (121)
$=$ $\frac{\pi}{2}[(2-m)\sum_{n=0}^{\infty}\frac{(\frac{1}{2})_{n}(\frac{1}{2})_{n}}{n!n!}m^{n}-2\sum_{n=0}^{\infty}\frac{(-\frac{1}{2})_{n}(\frac{1}{2})_{n}}{n!n!}m^{n}]$
$=$ $\frac{\pi}{2}\sum_{n=0}^{\infty}(2n(\frac{1}{2})_{n-1}-m(\frac{1}{2})_{n})\frac{(\frac{1}{2})_{n}}{n!n!}m^{n},$
where $(x)_{n}$ is Pochhammer’s symbol
$(x)_{n}=\left\{\begin{matrix}x(x+1)(x+2)\cdots(x+n-1)\ \ if\ n=1,2,3,...\\ 1\ \ if\ n=0\end{matrix}\right.$.
When $n\geq 1$,
$(2n(\frac{1}{2})_{n-1}-m(\frac{1}{2})_{n})\frac{(\frac{1}{2})_{n}}{n!n!}m^{n}>0,$
but
$(2n(\frac{1}{2})_{n-1}-m(\frac{1}{2})_{n})\frac{(\frac{1}{2})_{n}}{n!n!}m^{n}=-m$
at $n=0$. We sum (121) from $n=0$ to $3$,
$\sum_{n=0}^{3}(2n(\frac{1}{2})_{n-1}-m(\frac{1}{2})_{n})\frac{(\frac{1}{2})_{n}}{n!n!}m^{n}$
$=\frac{1}{8}m^{2}+\frac{7}{64}m^{3}+\frac{25}{256}m^{4}>0$
and therefore,
$(2-m)K(m)-2E(m)>0,\;\mbox{when}\;0\leq m<1.$
While if $m<0$, using imaginary-modulus transformation of [27, equation 19.7.5] one can transform the problem equivalent to the one with $0\leq m<1$. Specifically,
$(2-m)K(m)-2E(m)$ (122)
$=$ $\frac{2-m}{\sqrt{1-m}}K(\frac{-m}{1-m})-2\sqrt{1-m}E(\frac{-m}{1-m})$
$=$ $\frac{1}{1-m^{\prime}}[(2-m^{\prime})K(m^{\prime})-2E(m^{\prime})]>0$
where
$0\leq m^{\prime}=\frac{-m}{1-m}<1.$
Hence, (117) is Schur-convex, and thus has global minimum at
$S_{ii}=\Lambda\Lambda^{T}=\frac{\pi\tilde{\sigma}_{x}^{2}}{4},\quad i=1,2,$
and thus the minimum value of (22) when $d=2$ is $\tilde{\sigma}_{x}\sqrt{1-\pi/4}$.
Appendix F Estimate of the data CDF $\hat{F}_{\mu}(x)$ in (71)
we use kernel density estimation to estimate the data PDF
$\frac{1}{|\mathcal{M}|}\sum^{|\mathcal{M}|}_{i=1}\frac{1}{h}k_{i}\left(\frac{x-x_{i}}{h}\right),$
where the kernel $k_{i}(.)$ can be selected as Gaussian or the one easy to perform CDF inverse. The idea of kernel density estimation is a piecewise density estimation and it converges to the PDF when $|\mathcal{M}|$ is large [30]. To be specific, we first rearrange the data in set $\mathcal{M}$ such that
$x_{1}\leq x_{2}\leq\ldots\leq x_{|\mathcal{M}|}.$
Then if $k_{i}(x)$ is Epanechnikov which is $(3/4)(1-x^{2})$ if $x_{i}-h\leq x\leq x_{i}+h$ and 0 else, and by integration the total probability
$\int^{\infty}_{-\infty}\frac{1}{|\mathcal{M}|}\sum^{|\mathcal{M}|}_{i=1}\frac{1}{h}k_{i}\left(\frac{x-x_{i}}{h}\right)dx$
$=\frac{1}{|\mathcal{M}|}\sum_{i=1}^{|\mathcal{M}|}\int_{-1}^{1}\frac{3}{4}(1-y_{i}^{2})dy_{i}$
$=\frac{1}{|\mathcal{M}|}\sum_{i=1}^{|\mathcal{M}|}\frac{3}{4}\left[y_{i}-\frac{y_{i}^{3}}{3}\right]_{-1}^{1}=\frac{1}{|\mathcal{M}|}\sum_{i=1}^{|\mathcal{M}|}1=1$
where the change of variable is $y_{i}=\frac{x-x_{i}}{h},dx=hdy_{i}$. Then the CDF estimation is
$\hat{F}_{\mu}(x)=\int^{x}_{-\infty}\frac{1}{|\mathcal{M}|}\sum^{|\mathcal{M}|}_{i=1}\frac{1}{h}k_{i}\left(\frac{x^{\prime}-x_{i}}{h}\right)dx^{\prime}$
$=\frac{1}{|\mathcal{M}|}\left(\sum^{L^{x}}_{i=1}1+\int_{-1}^{(x-x_{L^{x}+1})/h}\frac{3}{4}(1-y_{L}^{2})dy_{L}\right)$
$=\frac{1}{|\mathcal{M}|}\left(L^{x}+\frac{3}{4}\left(\frac{x-x_{L^{x}+1}}{h}-\frac{(x-x_{L^{x}+1})^{3}}{3h^{3}}+1-\frac{1}{3}\right)\right)$
$=\frac{1}{|\mathcal{M}|}\left(L^{x}+\frac{1}{4}\left(2+3\left(\frac{x-x_{L^{x}+1}}{h}\right)-\left(\frac{x-x_{L^{x}+1}}{h}\right)^{3}\right)\right)$
if $x_{1}-h<x<x_{|\mathcal{M}|}+h$, where $L^{x}+1$ is the index of unique data point $x_{L^{x}+1}$ such that
$x_{L^{x}+1}-h<x<x_{L^{x}+1}+h.$
If $x\geq x_{|\mathcal{M}|}+h$ then $\hat{F}_{\mu}(x)=1$ and if $x\leq x_{1}-h$ then $\hat{F}_{\mu}(x)=0$. Also the inverse CDF
$\hat{F}^{-1}_{\mu}(p)=2\sin\left(\frac{1}{3}\arcsin\big{(}2\left(|\mathcal{M}|p-L^{p}\right)-1\big{)}\right)h+x_{L^{p}+1}$
where
$\frac{L^{p}}{|\mathcal{M}|}<p<\frac{L^{p}+1}{|\mathcal{M}|}.$